### v1

# Log-majorizations between quasi-geometric type means for matrices

Abstract
In this paper, for $\alpha\in(0,\infty)\setminus\{1\}$, $p>0$ and positive semidefinite matrices $A$ and $B$, we consider the quasi-extension $\mathcal{M}_{\alpha,p}(A,B):=\mathcal{M}_{\alpha}(A^{p},B^{p})^{1/p}$ of several $\alpha$-weighted geometric type matrix means $\mathcal{M}_{\alpha}(A,B)$ such as the $\alpha$-weighted geometric mean in Kubo–Ando’s sense, the Rényi mean, etc. The log-majorization $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ is examined for pairs $(\mathcal{M},\mathcal{N})$ of those $\alpha$-weighted geometric type means. The joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ is also discussed based on theory of quantum divergences.
2020 Mathematics Subject Classification: 15A45, 47A64
Keywords and phrases: $\alpha$-weighted geometric mean, $\alpha$-weighted spectral geometric mean, Rényi mean, Log-Euclidean mean, log-majorization, quantum channel, maximal $\alpha$-Rényi divergence, measured $\alpha$-Rényi divergence, $\alpha$-$z$-Rényi divergence
1 Introduction
Several types of majorizations have been well developed for the eigenvalues and the singular values of matrices, as is fully described in the book [43] (see also [2, 10, 29] for example), giving rise to powerful tools in deriving various matrix norm and trace inequalities. Among others, the notion of multiplicative type of majorization called log-majorization (see Section 2 for definition) has played an important role in matrix analysis, mathematical physics, quantum information, etc. For instance, for positive semidefinite matrices $A,B\geq 0$, Araki’s log-majorization [4] of Golden–Thompson type is
$(A^{p/2}B^{p}A^{p/2})^{1/p}\prec_{\log}(A^{q/2}B^{q}A^{q/2})^{1/q},\qquad 0<p\leq q,$ (1.1)
which is a stronger version of Araki–Lieb–Thirring trace inequality. Ando–Hiai’s one [3] of complementary Golden–Thompson type is
$(A^{p}\#_{\alpha}B^{p})^{1/p}\prec_{\log}(A^{q}\#_{\alpha}B^{q})^{1/q},\qquad 0<q\leq p,$ (1.2)
where $A\#_{\alpha}B$ is the $\alpha$-weighted geometric mean of $A,B$. On the other hand, for positive definite matrices $A,B>0$, Kian–Seo [39] obtained the counterpart of (1.2) for $\alpha\in[-1,0)\cup(1,2]$ as
$(A^{p}\#_{\alpha}B^{p})^{1/p}\prec_{\log}(A^{q}\#_{\alpha}B^{q})^{1/q},\qquad 0<p\leq q,$ (1.3)
where the definition of $A\#_{\alpha}B:=A^{1/2}(A^{-1/2}BA^{-1/2})^{\alpha}A^{1/2}$ for $A,B>0$ extends to all $\alpha\in\mathbb{R}$.
In our previous paper [33] we considered the quasi-arithmetic matrix mean $\mathcal{A}_{\alpha,p}(A,B):=\{(1-\alpha)A^{p}+\alpha B^{p}\}^{1/p}$ and several quasi-geometric type matrix means such as the quasi $\alpha$-weighted geometric mean $G_{\alpha,p}(A,B):=(A^{p}\#_{\alpha}B^{p})^{1/p}$, the quasi versions $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ of two different $\alpha$-weighted spectral geometric means, the Rényi mean $R_{\alpha,p}$, and the Log-Euclidean mean $LE_{\alpha}$ for $0<\alpha<1$ and $p>0$ (see Section 2 for the precise definitions of these). In [33] we examined various quasi-arithmetic-geometric inequalities between each of those quasi-geometric means and $\mathcal{A}_{\alpha,q}$ with respect to different matrix orderings. In the present paper we continue to consider the above mentioned quasi-geometric type means for all $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$. In fact, (1.1) is rewritten as $R_{\alpha,p}\prec_{\log}R_{\alpha,q}$, and (1.2) and (1.3) are nothing but $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$. Furthermore, the log-majorizations $G_{\alpha,p}\prec_{\log}R_{\alpha,q}$ and $R_{\alpha,q}\prec_{\log}G_{\alpha,p}$ were characterized in [31], and $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ was addressed in a recent paper [24]. In this paper we examine the log-majorization $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for other pairs $(\mathcal{M},\mathcal{N})$ from $R$, $G$, SGSG, $\widetilde{S}G$ and $LE$. Our goal is to hopefully obtain the necessary and sufficient condition on $p,q,\alpha$ under which $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ holds for all $A,B>0$, though we have not succeeded it for all cases. To do so, we need to address the sufficiency condition and the necessity condition separately. For the sufficiency part, since $\det\mathcal{M}_{\alpha,p}(A,B)=\det\mathcal{N}_{\alpha,q}(A,B)=(\det A)^{1-\alpha}(\det B)^{\alpha}$, the standard technique using anti-symmetric tensor powers is utilized so that the problem boils down to prove that
$\mathcal{N}_{\alpha,q}(A,B)\leq I\implies\mathcal{M}_{\alpha,p}(A,B)\leq I.$
The necessity part is quite computation-oriented. Similarly to the previous paper [33], we make computations for specific pairs of $2\times 2$ positive definite matrices to restrict the possible range of $p,q,\alpha$ for which $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ holds.
The joint concavity/convexity of matrix trace functions in two (or more) variables has been one of major subjects in matrix analysis since Lieb’s seminal paper [42] in 1973. Lieb’s concavity [42] and Ando’s convexity [1] are especially famous, and the joint concavity/convexity of the trace functions of the forms $\mathrm{Tr}(A^{p}+B^{q})^{s}$ and $\mathrm{Tr}(A^{p/2}B^{q}A^{p/2})^{s}$ was developed in, e.g., [15, 30]. The problem of fully characterizing the joint concavity/convexity of the above latter trace function became very important, because it settles the question of the monotonicity property under CPTP maps (or the data-processing inequality) of the $\alpha$-$z$-Rényi divergence [5], as explained in detail in [14]. Then the problem was finally solved by Zhang [53] in 2020, who characterized the joint concavity/convexity of a more general trace function $(A,B)\mapsto\mathrm{Tr}(\mathrm{Tr}(A^{p/2}X^{*}B^{q}XA^{p/2})^{s}$. Also, the joint concavity of the trace function $\mathrm{Tr}(A^{p}\sigma B^{q})^{s}$ for Kubo–Ando’s operator means $\sigma$ was shown in [30]. The current status of the subject has been summarized in Carlen’s recent survey paper [12] and his new book [13]. Our second aim of this paper is to examine the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for our quasi-geometric type means $\mathcal{M}_{\alpha,p}$ and for $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$, while the case $\mathrm{Tr}\,R_{\alpha,p}$ is included in Zhang’s result in [53], and $\mathrm{Tr}\,G_{\alpha,p}$ for $0<\alpha<1$ is a special case in [30]. Our idea is to apply the relationship between the joint concavity/convexity and the monotonicity under CPTP maps for geometric type matrix functions of two variables (see Theorem 5.3) similarly to the situation in the above mentioned monotonicity question of the $\alpha$-$z$-divergence. Then as a necessary condition for joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ we have the following sandwiched inequalities (see Theorem 5.8)
$\mathrm{Tr}\,G_{\alpha,1}(A,B)\leq\mathrm{Tr}\,\mathcal{M}_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)\quad\mbox{for $0<\alpha<1$},$
$\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)\leq\mathrm{Tr}\,\mathcal{M}_{\alpha,p}(A,B)\leq\mathrm{Tr}\,G_{\alpha,1}(A,B)\quad\mbox{for $\alpha>1$}.$
These inequalities can be analyzed from the log-majorizations between $\mathcal{M}_{\alpha,p}$ and $G_{\alpha,1}$, $R_{\alpha,1/\alpha}$. But it is left open to find a sufficient condition for $\mathcal{M}_{\alpha,p}$ to be jointly concave/convex, which seems a difficult problem.
The structure of the paper is as follows. In Section 2 we review the definitions of the quasi-geometric matrix means $R_{\alpha,p}$, $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$ mentioned above as well as the notion of log-majorization. We add a few basic properties of them to those given in [33]. The main Section 3 is divided into four subsections. In those subsections we examine the log-majorizations $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for all pairs $(\mathcal{M},\mathcal{N})$, except for the already known cases, from $R,G,SG.\widetilde{S}G,LE$. In Section 4 we characterize the equality cases in the norm inequalities derived from the log-majorizations shown in Section 3 in terms of the commutativity of matrix variables $A,B$, expanding the former results in [28, 32]. In Section 5 we discuss the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for $\mathcal{M}=R,G,SG,\widetilde{S}G,LE$ though some cases are not new, based on theory of quantum divergences. Finally in Section 6 some concluding remarks and open problems are in order. The paper contains three appendices. Appendix A is the proof of a main part of Theorem 4.1. Appendix B is the proof of Theorem 5.3 and Appendix C is a supplement to Theorems 5.3 and 5.8.
2 Preliminaries
For each $n\in\mathbb{N}$ we write $\mathbb{M}_{n}$ for the $n\times n$ complex matrices. Let $\mathbb{M}_{n}^{+}$ and $\mathbb{M}_{n}^{++}$ be the positive semidefinite $n\times n$ matrices and the positive definite $n\times n$ matrices, respectively. We often write $A\geq 0$ for $A\in\mathbb{M}_{n}^{+}$ and $A>0$ for $A\in\mathbb{M}_{n}^{++}$ (for some $n\in\mathbb{N}$). The $n\times n$ identity matrix is denoted by $I_{n}$ or simply $I$. Let $\mathrm{Tr}$ be the usual trace on $\mathbb{M}_{n}$ and $\|X\|_{\infty}$ be the operator norm of $X\in\mathbb{M}_{n}$. For $A\geq 0$ we write $s(A)$ for the support projection of $A$. We write $A^{-1}$ for the generalized inverse of $A$, i.e., the inverse of $A$ under the restriction to the support of $A$. Moreover, for $r<0$ we define $A^{r}:=(A^{-1})^{-r}$ via the generalized inverse.
In this preliminary section we recall several examples of quasi matrix means and a few notions of matrix orders, in particular, log-majorization. We first enumerate the definitions of quasi-extensions of several binary matrix means for matrices. Let $0<\alpha<1$ and $p>0$, and let $A,B\in\mathbb{M}_{n}^{+}$.
(i) The quasi $\alpha$-weighted arithmetic mean is
$\mathcal{A}_{\alpha,p}(A,B):=(A^{p}\triangledown_{\alpha}B^{p})^{1/p}=((1-\alpha)A^{p}+\alpha B^{p})^{1/p},$
which is also called the ($\alpha$-weighted) matrix $p$-power mean.
(ii) The $\alpha$-weighted harmonic mean of $A,B>0$ is $A!_{\alpha}B:=((1-\alpha)A^{-1}+\alpha B^{-1})^{-1}$, extended to general $A,B\geq 0$ as $A!_{\alpha}B:=\lim_{\varepsilon\searrow 0}(A+\varepsilon I)!_{\alpha}(B+\varepsilon I)$. The quasi $\alpha$-weighted harmonic mean is
$\mathcal{H}_{\alpha,p}(A,B):=(A^{p}!_{\alpha}B^{p})^{1/p}.$
(iii) The $\alpha$-geometric mean of $A,B>0$ is
$A\#_{\alpha}B:=A^{1/2}(A^{-1/2}BA^{-1/2})^{\alpha}A^{1/2},$ (2.1)
which is extended to $A,B\geq 0$ as $A\#_{\alpha}B:=\lim_{\varepsilon\searrow 0}(A+\varepsilon I)\#_{\alpha}(B+\varepsilon I)$. The quasi $\alpha$-weighted geometric mean is
$G_{\alpha,p}(A,B):=(A^{p}\#_{\alpha}B^{p})^{1/p}.$ (2.2)
(iv) The spectral geometric mean of $A,B>0$ due to Fiedler and Pták [20] is
$F(A,B):=(A^{-1}\#B)^{1/2}A(A^{-1}\#B)^{1/2},$
which was extended to the $\alpha$-weighted version in [41] as
$F_{\alpha}(A,B):=(A^{-1}\#B)^{\alpha}A(A^{-1}\#B)^{\alpha}.$ (2.3)
The $\alpha$-weighted spectral geometric mean has recently been studied in [40, 22, 24, 21] where $F_{\alpha}(A,B)$ is denoted by $A\natural_{\alpha}B$. The above definition of $F(A,B)$ is also meaningful for $A,B\geq 0$ with $s(A)\geq s(B)$ where $A^{-1}$ is the generalized inverse. The quasi $\alpha$-weighted spectral geometric mean of $A,B\geq 0$ with $s(A)\geq s(B)$ is
$SG_{\alpha,p}(A,B):=F_{\alpha}(A^{p},B^{p})^{1/p}.$ (2.4)
(v) Another weighted version of the spectral geometric mean of $A,B>0$ recently introduced in [16] is
$\widetilde{F}_{\alpha}(A,B):=(A^{-1}\#_{\alpha}B)^{1/2}A^{2(1-\alpha)}(A^{-1}\#_{\alpha}B)^{1/2},$ (2.5)
whose quasi-extension is
$\widetilde{S}G_{\alpha,p}(A,B):=\widetilde{F}_{\alpha}(A^{p},B^{p})^{1/p}.$ (2.6)
These $\widetilde{F}_{\alpha}(A,B)$ and $\widetilde{S}G_{\alpha,p}(A,B)$ are meaningful for $A,B\geq 0$ with $s(A)\geq s(B)$ as well.
(vi) There is one more familiar quasi matrix mean defined for all $A,B\geq 0$ by
$R_{\alpha,p}(A,B):=\bigl(A^{{1-\alpha\over 2}p}B^{\alpha p}A^{{1-\alpha\over 2}p}\bigr)^{1/p}.$ (2.7)
This is called the Rényi mean in [17] because $\mathrm{Tr}\,R_{\alpha,p}$ appears as the main component in the definition of a certain quantum Rényi divergence; see Section 5.2 for more details.
(vii) The Log-Euclidean mean of $A,B>0$ is
$LE_{\alpha}(A,B):=\exp((1-\alpha)\log A+\alpha\log B),$
which is extended to general $A,B\geq 0$ as
$LE_{\alpha}(A,B):=P_{0}\exp\{(1-\alpha)P_{0}(\log A)P_{0}+\alpha P_{0}(\log B)P_{0}\},$ (2.8)
where $P_{0}:=s(A)\wedge s(B)$. There is no quasi-extension of $LE_{\alpha}$ because of $LE_{\alpha}(A^{p},B^{p})^{1/p}=LE_{\alpha}(A,B)$ for all $p>0$.
In this paper we will consider quasi-geometric type matrix means $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$, $R_{\alpha,p}$ and $LE_{\alpha}$ for not only $0<\alpha<1$ but also $\alpha>1$. The above definitions in (2.1)–(2.8) are all available even for any $\alpha>0$, $p>0$ and for all $A,B\geq 0$ with $s(A)\geq s(B)$ under conventions of $A^{-1}$ and $A^{r}$ for $r<0$ mentioned in the beginning of this section. To be precise, we here fix the domains of those quasi-geometric type means as follows: The domain of $G_{\alpha,p},R_{\alpha,p},LE_{\alpha}$ for $0<\alpha<1$ is $\bigsqcup_{n\geq 1}(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})$. The domain of $G_{\alpha,p},R_{\alpha,p},LE_{\alpha}$ for $\alpha>1$ and that of $SG_{\alpha,p},\widetilde{S}G_{\alpha,p}$ for all $\alpha\in(0,\infty)\setminus\{1\}$ are
$\bigsqcup_{n\geq 1}\{(A,B)\in\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}:s(A)\geq s(B)\}.$
In this way, although our quasi-geometric type means for $\alpha>1$ are a bit less meaningful as matrix means than those for $0<\alpha<1$, we consider those for all $\alpha\in(0,\infty)\setminus\{1\}$ in this paper.
Several basic facts on the quasi matrix means defined in (i)–(vii) have been collected, though restricted to the case $0<\alpha<1$, in Section 2.1 of our previous paper [33]. A few of those are supplemented by the next proposition and theorem including the case $\alpha>1$.
Proposition 2.1.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$. Let $\mathcal{M}_{\alpha,p}$ be any of $G_{\alpha,p}$, $R_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$.
  • (1)
    $\mathcal{M}_{\alpha,p}(A^{-1},B^{-1})=\mathcal{M}_{\alpha,p}(A,B)^{-1}$ for all $A,B>0$.
  • (2)
    For every $(A,B)$ in the domain of $\mathcal{M}_{\alpha,p}$ we have $\mathcal{M}_{\alpha,p}(A,B)=\lim_{\varepsilon\searrow 0}\mathcal{M}_{\alpha,p}(A+\varepsilon I,B+\varepsilon I)$.
Proof.
(1) is easily verified by definition of each $\mathcal{M}_{\alpha,p}$.
(2) For $0<\alpha<1$ the assertion was shown in [33, Proposition 2.2]. Let $\alpha>1$ and $A,B\geq 0$ with $s(A)\geq s(B)$. Then the proof of [33, Proposition 2.2] for $SG_{\alpha,p}$ and $\widetilde{S}G_{\alpha,p}$ with $0<\alpha<1$ can work for any of $G_{\alpha,p}$, $R_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ with $\alpha>1$ as well. The proof for $LE_{\alpha}$ is similar to that of [36, Lemma 4.1], while we here give a short proof using [33, Lemma A.3]. Let $P_{0}:=s(B)$ and write for $0<\varepsilon<1$,
$Z(\varepsilon):=(1-\alpha)\log(A+\varepsilon I)+\alpha\log(B+\varepsilon I)=\begin{bmatrix}Z_{0}(\varepsilon)&Z_{2}(\varepsilon)\\ Z_{2}^{*}(\varepsilon)&Z_{1}(\varepsilon)\end{bmatrix}$,
where $Z_{0}(\varepsilon):=P_{0}Z(\varepsilon)P_{0}$, $Z_{1}(\varepsilon):=P_{0}^{\perp}Z(\varepsilon)P_{0}^{\perp}$ and $Z_{2}(\varepsilon):=P_{0}Z(\varepsilon)P_{0}^{\perp}$. Since
$Z_{0}(\varepsilon)$ $=(1-\alpha)P_{0}(\log(A+\varepsilon I))P_{0}+\alpha P_{0}(\log(B+\varepsilon I))P_{0},$
$Z_{1}(\varepsilon)$ $=(1-\alpha)P_{0}^{\perp}(\log(A+\varepsilon I))P_{0}^{\perp}+\alpha(\log\varepsilon)P_{0}^{\perp},$
$Z_{2}(\varepsilon)$ $=(1-\alpha)P_{0}(\log(A+\varepsilon I))P_{0}^{\perp},$
it is easy to see that
$Z_{0}(\varepsilon)\to Z_{0}:=(1-\alpha)P_{0}(\log A)P_{0}+\alpha P_{0}(\log B)P_{0}\quad\mbox{as $\varepsilon\searrow 0$},$
${1\over-\log\varepsilon}\,Z_{1}(\varepsilon)={\alpha-1\over\log\varepsilon}\,P_{0}^{\perp}(\log(A+\varepsilon I))P_{0}^{\perp}-\alpha P_{0}^{\perp}\to-\alpha P_{0}^{\perp}\quad\mbox{as $\varepsilon\searrow 0$},$
$\sup\{\|Z_{2}(\varepsilon)\|_{\infty}:0<\varepsilon<1\}<\infty.$
Hence we can apply [33, Lemma A.3] to $Z(\varepsilon)$ with parameter $p:={1\over-\log\varepsilon}$ ($\searrow 0$ as $\varepsilon\searrow 0$) to obtain
$LE_{\alpha}(A+\varepsilon I,B+\varepsilon I)=e^{Z(\varepsilon)}$
$\quad\to P_{0}e^{Z_{0}}=P_{0}\exp\{(1-\alpha)P_{0}(\log A)P_{0}+\alpha P_{0}(\log B)P_{0}\}=LE_{\alpha}(A,B),$
as desired. ∎
Theorem 2.2.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$. Let $\mathcal{M}_{\alpha,p}$ be any of $G_{\alpha,p}$, $R_{\alpha,p}$, $SG_{\alpha,p}$ and $\widetilde{S}G_{\alpha,p}$. For every $(A,B)$ in the domain of $\mathcal{M}_{\alpha,p}$ we have $LE_{\alpha}(A,B)=\lim_{p\searrow 0}\mathcal{M}_{\alpha,p}(A,B)$.
To prove the theorem we give a lemma. Note that for $0<\alpha<1$ this is contained in [33, (A.25)] where $A,B\geq 0$ are general and $P_{0}:=s(A)\wedge s(B)$.
Lemma 2.3.
Let $\alpha,p$ be as in Theorem 2.2. For every $A,B\geq 0$ with $s(A)\geq s(B)$ we have
$A^{p}\#_{\alpha}B^{p}=P_{0}+p\{(1-\alpha)P_{0}(\log A)P_{0}+\alpha P_{0}(\log B)P_{0}\}+o(p)\quad\mbox{as $p\searrow 0$},$
where $P_{0}:=s(B)$.
Proof.
Let $L:=P_{0}(-\log A)P_{0}+P_{0}(\log B)P_{0}$ and set $Y(p):=(A^{-p/2}B^{p}A^{-p/2})^{1/p}-P_{0}e^{L}P_{0}$. Then by [33, (A.2)] we have $Y(p)\to 0$ as $p\searrow 0$. We write
$A^{-p/2}B^{p}A^{-p/2}$ $=P_{0}(e^{L}+Y(p))^{p}P_{0}=P_{0}\bigl\{\exp\bigl[p\log(e^{L}+Y(p))\bigr]\bigr\}P_{0}$ (2.9)
$=P_{0}\bigl\{I+p\log(e^{L}+Y(p))+o(p)\bigr\}P_{0}=P_{0}+pL+o(p)\quad\mbox{as $p\searrow 0$},$
where the last equality follows since Taylor’s theorem (see, e.g., [29, Theorem 2.3.1]) gives
$\log(e^{L}+Y(p))=L+D(\log x)(e^{L})(Y(p))+o(1)=L+o(1)$
with the Fréchet derivative $D(\log x)(e^{L})$ of the functional calculus by $\log x$ ($x>0$) at $e^{L}$ (see [29, p. 159]). Hence the Taylor expansion implies that $(A^{-p/2}B^{p}A^{-p/2})^{\alpha}=P_{0}+\alpha pL+o(p)$ as $p\searrow 0$. Moreover, note that $A^{p}\#_{\alpha}B^{p}=P_{0}(A^{p}\#_{\alpha}B^{p})P_{0}$. This is obvious for $0<\alpha<1$ (even for general $A,B\geq 0$). For $\alpha>1$ this can be seen from
$A^{p}\#_{\alpha}B^{p}$ $=A^{p/2}(A^{-p/2}B^{p}A^{-p/2})^{\alpha}A^{p/2}$
$=B^{p}A^{-p/2}(A^{-p/2}B^{p}A^{-p/2})^{\alpha-1}A^{p/2}=A^{p/2}(A^{-p/2}B^{p}A^{-p/2})^{\alpha-1}A^{-p/2}B^{p}.$
Therefore, we have
$A^{p}\#_{\alpha}B^{p}$ $=P_{0}\bigl\{A^{p/2}(A^{-p/2}B^{p}A^{-p/2})^{\alpha}A^{p/2}\bigr\}P_{0}$
$=P_{0}\Bigl(s(A)+{p\over 2}\,s(A)\log A+o(p)\Bigr)(P_{0}+pL+o(p))\Bigl(s(A)+{p\over 2}\,s(A)\log A+o(p)\Bigr)P_{0}$
$=P_{0}+pP_{0}(\log A)P_{0}+\alpha pL+o(p)$
$=P_{0}+p\{(1-\alpha)P_{0}(\log A)P_{0}+\alpha P_{0}(\log B)P_{0}\}+o(p)\quad\mbox{as $p\searrow 0$},$
as asserted. ∎
Proof of Theorem 2.2.
For $0<\alpha<1$ the assertion was shown in [33, Theorem 2.3]. For $\alpha>1$ let $A,B\geq 0$ with $s(A)\geq s(B)$. For $G_{\alpha,p}$ Lemma 2.3 gives $G_{\alpha,p}(A,B)^{p}=P_{0}+pK+o(p)$, where $K:=(1-\alpha)P_{0}(\log A)P_{0}+\alpha P_{0}(\log B)P_{0}$. This implies that with $\mathcal{H}_{0}$ being the range of $P_{0}$,
$\log G_{\alpha,p}(A,B)\big|_{\mathcal{H}_{0}}={1\over p}\log\{P_{0}+pK+o(p)\}\big|_{\mathcal{H}_{0}}=(K+o(1))|_{\mathcal{H}_{0}},$
showing the result for $G_{\alpha,p}$. For $R_{\alpha,p}$ apply (2.9) to $A^{\alpha-1}$ and $B^{\alpha}$ in place of $A,B$ to have $A^{{1-\alpha\over 2}p}B^{\alpha p}A^{{1-\alpha\over 2}p}=P_{0}+pK+o(p)$. This implies that $\log R_{\alpha,p}(A,B)\big|_{\mathcal{H}_{0}}=(K+o(1))|_{\mathcal{H}_{0}}$, showing the result for $R_{\alpha,p}$. The proof for $SG_{\alpha,p}$ is similar to that of [33, Theorem 2.3] for $0<\alpha<1$. Finally, for $\widetilde{S}G_{\alpha,p}$ by Lemma 2.3 we have $A^{-p}\#_{\alpha}B^{p}=P_{0}+p\widetilde{K}+o(p)$, where $\widetilde{K}:=(1-\alpha)P_{0}(-\log A)P_{0}+\alpha P_{0}(\log B)P_{0}$. Therefore,
$(A^{-p}\#_{\alpha}B^{p})^{1/2}A^{2(1-\alpha)p}(A^{-p}\#_{\alpha}B^{p})^{1/2}$
$\qquad=\Bigl(P_{0}+{p\over 2}\,\widetilde{K}+o(p)\Bigr)(s(A)+2(1-\alpha)ps(A)\log A+o(p))\Bigl(P_{0}+{p\over 2}\,\widetilde{K}+o(p)\Bigr)$
$\qquad=P_{0}+2(1-\alpha)pP_{0}(\log A)P_{0}+p\widetilde{K}+o(p)=P_{0}+pK+o(p),$
which shows the result for $\widetilde{S}G_{\alpha,p}$. ∎
In our previous paper [33] we discussed arithmetic-geometric type inequalities for quasi-geometric type matrix means in several different matrix orderings varying from the strongest Loewner order to the weakest order determined by trace inequality. But in the present paper we are mostly concerned with log-majorizations for quasi-geometric type matrix means stated in (iii)–(vii) above.
Let $X,Y\in\mathbb{M}_{n}^{+}$. The most standard and the strongest order between $X,Y$ is the the Loewner order $X\leq Y$, i.e., $Y-X\geq 0$. Let $\lambda(X)=(\lambda_{1}(X),\dots,\lambda_{n}(X))$ be the eigenvalues of $X$ in decreasing order with multiplicities. The entrywise eigenvalue order denoted as $X\leq_{\lambda}Y$ is defined if $\lambda_{i}(X)\leq\lambda_{i}(Y)$ for each $i=1,\dots,n$. The weak log-majorization $X\prec_{w\log}Y$ means that
$\prod_{i=1}^{k}\lambda_{i}(X)\leq\prod_{i=1}^{k}\lambda_{i}(Y),\qquad 1\leq k\leq n,$
and the log-majorization $X\prec_{\log}Y$ means that $X\prec_{w\log}Y$ and $\prod_{i=1}^{n}\lambda_{i}(X)=\prod_{i=1}^{n}\lambda_{i}(Y)$, i.e., $\det X=\det Y$. Details on (weak) log-majorization are found in [2], [10, Chap. II] and [43]. Some basic properties of several matrix orderings including $\leq,\leq_{\lambda},\prec_{(w)\log}$ mentioned above are also found in [33, Sec. 2.2].
Concerning the quasi-geometric type matrix means in (iii)–(vii) above, some general facts are in order.
Remark 2.4.
(1) Let $(\mathcal{M},\mathcal{N})$ be any pair from $G,SG,\widetilde{S}G,R,LE$, and let $\alpha,\beta\in(0,\infty)\setminus\{1\}$ and $p,q>0$. When $\alpha\neq\beta$, $\mathcal{M}_{\alpha,p}$ and $\mathcal{N}_{\beta,q}$ are not definitively comparable even for positive scalar variables, as explained in [33, Remark 2.7(1)]. So we only compare between $\mathcal{M}_{\alpha,p}$ and $\mathcal{N}_{\alpha,q}$ under the same weight parameter $\alpha$.
(2) Let $\mathcal{M}_{\alpha,p}$ and $\mathcal{N}_{\alpha,q}$ be as in (1), and let $A,B>0$. Since
$\det\mathcal{M}_{\alpha,p}(A,B)=\det\mathcal{N}_{\alpha,q}(A,B)=(\det A)^{1-\alpha}(\det B)^{\alpha}$
independently of $p,q>0$, we have $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}$ as long as $\mathcal{M}_{\alpha,p}(A,B)\prec_{w\log}\mathcal{N}_{\alpha,q}(A,B)$. Furthermore, if $\mathcal{M}_{\alpha,p}(A,B)\leq_{\lambda}\mathcal{N}_{\alpha,q}(A,B)$, then we must have $\lambda(\mathcal{M}_{\alpha,p}(A,B))=\lambda(\mathcal{N}_{\alpha,q}(A,B))$. These facts are reasons why we consider only the log-majorization $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ in this paper.
(3) For any $p>0$ and $A,B\geq 0$ we note that
$\lambda(R_{1/2,2p}(A,B))=\lambda^{1/p}(A^{p/2}B^{p}A^{p/2})=\lambda^{1/p}(B^{p/2}A^{p}B^{p/2})=\lambda(SG_{1/2,p}(A,B))$
thanks to [20, Theorem 3.2, Item 8] and that $SG_{1/2,p}(A,B)=\widetilde{S}G_{1/2,p}(A,B)$ by definitions. Moreover, for $A,B\geq 0$ with $s(A)\geq s(B)$ we have
$\lambda(R_{2,p}(A,B))=\lambda^{1/p}(A^{-p/2}B^{2p}A^{-p/2})=\lambda^{1/p}(B^{p}A^{-p}B^{p})=\lambda(G_{2,p}(A,B)).$
Hence the three pairs $(R_{1/2,2p},SG_{1/2,p})$, $(R_{2,p},G_{2,p})$ and $(SG_{1/2,p},\widetilde{S}G_{1/2,p})$ are rather trivial and exceptional cases in our study.
(4) For any pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ except for the three cases mentioned in (3), note that there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\mathcal{M}_{\alpha,p}(A,B)\not\leq_{\lambda}\mathcal{N}_{\alpha,q}(A,B)$. Indeed, assume that this is not the case. Then by (2) we must have $\lambda(\mathcal{M}_{\alpha,p}(A,B))=\lambda(\mathcal{N}_{\alpha,q}(A,B))$ for all $A,B\in\mathbb{M}_{2}^{++}$. But this fails to hold except for the above three cases and the trivial case of $\mathcal{M}=\mathcal{N}$ and $p=q$.
For $x,y>0$ and $\theta\in\mathbb{R}$ we define $2\times 2$ positive definite matrices by
$A_{0}:=\begin{bmatrix}1&0\\ 0&x\end{bmatrix},\qquad B_{\theta}:=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}1&0\\ 0&y\end{bmatrix}\begin{bmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{bmatrix}$. (2.10)
These $A_{0},B_{\theta}$ were repeatedly utilized in [33] and are also useful in the present paper. The next lemma is used to verify the assertion of Remark 2.4(4), and it will be also used in several places in Section 3.
Lemma 2.5.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$. Let $A_{0},B_{\theta}\in\mathbb{M}_{2}^{++}$ be given in (2.10) with $y=x\in(0,1)$. Then we have
$\lambda_{1}(R_{\alpha,p}(A_{0},B_{\theta}))$ $=1+\theta^{2}\,{-1-x^{p}+x^{\alpha p}+x^{(1-\alpha)p}\over p(1-x^{p})}+o(\theta^{2}),$ (2.11)
$\lambda_{1}(G_{\alpha,p}(A_{0},B_{\theta}))$ $=1+\theta^{2}\,{\alpha(1-\alpha)\over 2p}(x^{p}-x^{-p})+o(\theta^{2}),$ (2.12)
$\lambda_{1}(SG_{\alpha,p}(A_{0},B_{\theta}))$ $=1-\theta^{2}\,{2\alpha(1-\alpha)\over p}\cdot{1-x^{p}\over 1+x^{p}}+o(\theta^{2}),$ (2.13)
$\lambda_{1}(\widetilde{S}G_{\alpha,p}(A_{0},B_{\theta}))$ $=1-{\theta^{2}\over p}\biggl\{\alpha\,{1-x^{p}\over 1+x^{p}}+{x^{p}+x^{2p}-x^{(2\alpha+1)p}-x^{2(1-\alpha)p}\over(1-x^{p})(1+x^{p})^{2}}\biggr\}+o(\theta^{2}),$ (2.14)
$\lambda_{1}(LE_{\alpha}(A_{0},B_{\theta}))$ $=1+\theta^{2}\alpha(1-\alpha)\log x+o(\theta^{2}).$ (2.15)
Proof.
We can easily compute all the expressions in the lemma by applying [33, Lemma 3.6] to the relevant situations discussed in [33]. See [33, Lemmas 3.4, 3.5 and 4.17] (with $y=x$) for (2.15), (2.11) and (2.12) respectively, and see the proofs of [33, Theorems 4.29 and 4.37] for (2.13) and (2.14) respectively, while details of computations are left to the reader. Also we note that all the relevant expressions in [33] for $0<\alpha<1$ remain valid for all $\alpha\in(0,\infty)\setminus\{1\}$. ∎
To verify Remark 2.4(4) more explicitly, we here include brief discussions for completeness. Consider $A_{0},B_{\theta}\in\mathbb{M}_{2}^{++}$ as in Lemma 2.5. First let $(\mathcal{M}_{\alpha,p},\mathcal{M}_{\alpha,q})$ be be given for $\mathcal{M}\in\{R,G,SG,\widetilde{S}G\}$. Assume that $\lambda_{1}(\mathcal{M}_{\alpha,p}(A_{0},B_{\theta}))=\lambda_{1}(\mathcal{M}_{\alpha,q}(A_{0},B_{\theta}))$ for all $x\in(0,1)$, which implies from (2.11)–(2.15) that ${-1-x^{p}+x^{\alpha p}+x^{(1-\alpha)p}\over p(1-x^{p})}={-1-x^{q}+x^{\alpha q}+x^{(1-\alpha)q}\over q(1-x^{q})}$ etc. Letting $x\searrow 0$ gives $p=q$ in each case. Next let $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ be given for $\mathcal{M}\neq\mathcal{N}$, and assume that $\lambda_{1}(\mathcal{M}_{\alpha,p}(A_{0},B_{\theta}))=\lambda_{1}(\mathcal{N}_{\alpha,q}(A_{0},B_{\theta}))$ for all $x\in(0,1)$. When $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})=(R_{\alpha,p},G_{\alpha,q})$, by (2.11) and (2.12) we have
${-1-x^{p}+x^{\alpha p}+x^{(1-\alpha)p}\over p(1-x^{p})}={\alpha(1-\alpha)\over 2q}(x^{q}-x^{-q}),\qquad x\in(0,1).$ (2.16)
For $0<\alpha<1$, letting $x\searrow 0$ gives $1\over p$$\infty$, a contradiction. For $\alpha>1$, we have $(1-\alpha)p=-q$ and ${1\over p}={\alpha(\alpha-1)\over 2q}$ by comparing the leading terms as $x\searrow 0$ of both sides of (2.16). Hence $\alpha=2$ and $p=q$, which is a case excluded. For other pairs the discussions are more or less similar, so we omit the details.
3 Log-majorizations
Let us begin with surveying the log-majorizations known so far between quasi-geometric type means $R_{\alpha,p}$, $G_{\alpha,p}$ and $SG_{\alpha,p}$ in the following theorem.
Theorem 3.1.
Let $\alpha\in(0,\infty)\setminus\{1\}$, $p,q>0$ and $A,B>0$.
  • (a)
    $R_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for any $\alpha\in(0,\infty)\setminus\{1\}$ if $p\leq q$.
  • (b)
    $G_{\alpha,q}(A,B)\prec_{\log}G_{\alpha,p}(A,B)$ if $0<\alpha<1$ and $p\leq q$.
  • (c)
    $G_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,q}(A,B)$ if $1<\alpha\leq 2$ and $p\leq q$.
  • (d)
    $G_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,q}(A,B)$ if $\alpha\geq 2$ and $p/q\leq{\alpha\over 2(\alpha-1)}$.
  • (e)
    $G_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ either if $0<\alpha<1$ and $p,q>0$ are arbitrary, or if $\alpha>1$ and $p/q\leq\min\{\alpha/2,\alpha-1\}$.
  • (f)
    $R_{\alpha,q}(A,B)\prec_{\log}G_{\alpha,p}(A,B)$ if $\alpha>1$ and $p/q\geq\max\{\alpha/2,\alpha-1\}$.
  • (g)
    $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ if $0<\alpha<1$ and $p/q\geq\max\{\alpha,1-\alpha\}$.
Indeed, (a) and (b) are the log-majorizations due to Araki [4] and Ando and Hiai [3, Theorem 2.1]), respectively. In [39] the matrix perspective $G_{\beta}(A,B):=A^{1/2}(A^{-1/2}BA^{-1/2})^{\beta}A^{1/2}$ for $1\leq\beta<0$ was treated, where $G_{\beta}(A,B)$ is denoted by $A\natural_{\beta}B$. Since $G_{\beta}(A,B)=G_{1-\beta}(B,A)$ and $1<1-\beta\leq 2$, the log-majorization in (c) is equivalent to that shown by Kian and Seo [39, Theorem 3.1]. (d) was shown in [31, Corollary 5.2], and (e) and (f) were shown in [31, Proposition 5.1(a), (b)]. Finally, (g) is a rewriting of the log-majorization due to Gan and Tam [24, Theorem 3.7]. By Proposition 2.1(2) note that the log-majorizations in the $0<\alpha<1$ case of (a), (b) and (e) hold for all $A,B\geq 0$ and those in the $\alpha>1$ case of (a), (c), (d), (f) and (g) hold for $A,B\geq 0$ with $s(A)\geq s(B)$.
In this section we will consider log-majorizations $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for pairs $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ of quasi-geometric type matrix means other than the above known cases. Our discussions will be divided into four subsections.
3.1 $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ and $R_{\alpha,q}\prec_{\log}SG_{\alpha,p}$
First we recall that the log-majorizations in (e) and (f) of Theorem 3.1 were indeed presented in [31] in the ‘if and only if’ statement. For the convenience to make explicit comparison with the log-majorizations shown in Section 3 below, we include [31, Proposition 5.1] in its complete form in the following:
Theorem 3.2.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p,q>0$.
  • (1)
    The following conditions are equivalent:
    • (i)
      $G_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$;
    • (ii)
      $G_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$;
    • (iii)
      either $0<\alpha<1$ and $p,q>0$ are arbitrary, or $\alpha>1$ and $p/q\leq\min\{\alpha/2,\alpha-1\}$.
  • (2)
    The following conditions are equivalent:
    • (i)
      $R_{\alpha,q}(A,B)\prec_{\log}G_{\alpha,p}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$;
    • (ii)
      $R_{\alpha,q}(A,B)\prec_{\log}G_{\alpha,p}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$;
    • (iii)
      $\alpha>1$ and $p/q\geq\max\{\alpha/2,\alpha-1\}$.
When $0<\alpha<1$, the next theorem gives the necessary and sufficient condition on $p,q$ for $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ (resp., $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$) to hold for all $A,B\geq 0$ with $s(A)\geq s(B)$, thus strengthening the assertion in (g) above.
Theorem 3.3.
Let $0<\alpha<1$ and $p,q>0$.
  • (1)
    The following conditions are equivalent:
    • (i)
      $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$;
    • (ii)
      $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$;
    • (iii)
      $p/q\leq\min\{\alpha,1-\alpha\}$.
  • (2)
    The following conditions are equivalent:
    • (i)
      $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$;
    • (ii)
      $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$;
    • (iii)
      $p/q\geq\max\{\alpha,1-\alpha\}$.
Proof.
[…]
$(X^{\alpha}YX^{\alpha})^{r}\prec_{\log}Y^{{1-\alpha\over 2}r}(XYX)^{\alpha r}Y^{{1-\alpha\over 2}r},$
where $r:=q/p$. Conversely, for any $X,Y>0$, setting $A:=Y^{1/p}$ and $B:=(XYX)^{1/p}$ we have $Y=A^{p}$ and $X=A^{-p}\#B^{p}$. Furthermore, note that for $X_{1},X_{2}\in\mathbb{M}_{2}^{++}$ with $\det X_{1}=\det X_{2}$, $X_{1}\prec_{\log}X_{2}$ if and only if $\lambda_{1}(X_{1})\leq\lambda(X_{2})$. Hence we see that condition (ii) is equivalent to saying that
$\lambda_{1}^{r}(X^{\alpha}YX^{\alpha})\leq\lambda_{1}\bigl(Y^{{1-\alpha\over 2}r}(XYX)^{\alpha r}Y^{{1-\alpha\over 2}r}\bigr)$ (3.1)
for all $X,Y\in\mathbb{M}_{2}^{++}$.
Now let $X:=A_{0}$ and $Y:=B_{\theta}$ in (2.10) for $x,y>0$ with $x^{2}y\neq 1$ and $x^{2\alpha}y\neq 1$, and argue as in the proof of [33, Theorem 4.23]. Since
$XYX=\begin{bmatrix}1-\theta^{2}(1-y)&\theta x(1-y)\\ \theta x(1-y)&x^{2}y+\theta^{2}x^{2}(1-y)\end{bmatrix}+o(\theta^{2})\quad\mbox{as $\theta\to 0$},$ (3.2)
one can apply [33, Example 3.2(1)] to compute
$(XYX)^{\alpha r}=\begin{bmatrix}1+\theta^{2}a_{11}&\theta a_{12}\\ \theta a_{12}&x^{2\alpha r}y^{\alpha r}+\theta^{2}a_{22}\end{bmatrix}+o(\theta^{2}),$
where
$\begin{cases}a_{11}:=-\alpha r(1-y)+{x^{2}(1-y)^{2}(\alpha r-1-\alpha rx^{2}y+x^{2\alpha r}y^{\alpha r})\over(1-x^{2}y)^{2}},\\ a_{12}:={x(1-y)(1-x^{2\alpha r}y^{\alpha r})\over 1-x^{2}y}.\end{cases}$
(The explicit form of $a_{22}$ is unnecessary below.) Since
$Y^{{1-\alpha\over 2}r}=\begin{bmatrix}1-\theta^{2}\bigl(1-y^{{1-\alpha\over 2}r}\bigr)&\theta\bigl(1-y^{{1-\alpha\over 2}r}\bigr)\\ \theta\bigl(1-y^{{1-\alpha\over 2}r}\bigr)&y^{{1-\alpha\over 2}r}+\theta^{2}\bigl(1-y^{{1-\alpha\over 2}r}\bigr)\end{bmatrix}+o(\theta^{2}),$
one further computes
$Y^{{1-\alpha\over 2}r}(XYX)^{\alpha r}Y^{{1-\alpha\over 2}r}=\begin{bmatrix}1+\theta^{2}b_{11}&\theta b_{12}\\ \theta b_{12}&x^{2\alpha r}y^{r}+\theta^{2}b_{22}\end{bmatrix}+o(\theta^{2}),$ (3.3)
where
$\begin{cases}b_{11}:=-2\bigl(1-y^{{1-\alpha\over 2}r}\bigr)+x^{2\alpha r}y^{\alpha r}\bigl(1-y^{{1-\alpha\over 2}r}\bigr)^{2}+a_{11}+2\bigl(1-y^{{1-\alpha\over 2}r}\bigr)a_{12},\\ b_{12}:=1-y^{{1-\alpha\over 2}r}+x^{2\alpha r}\bigl(y^{{1+\alpha\over 2}r}-y^{r}\bigr)+y^{{1-\alpha\over 2}r}a_{12}.\end{cases}$
On the other hand, one has
$X^{\alpha}YX^{\alpha}=\begin{bmatrix}1-\theta^{2}(1-y)&\theta x^{\alpha}(1-y)\\ \theta x^{\alpha}(1-y)&x^{2\alpha}y+\theta^{2}x^{2\alpha}(1-y)\end{bmatrix}+o(\theta^{2}),$ (3.4)
When $x^{2\alpha}y<1$, we apply [33, Lemma 3.6] to (3.4) and (3.3) to obtain
$\lambda_{1}^{r}(X^{\alpha}YX^{\alpha})$ $=1+\theta^{2}r\biggl(-1+y+{x^{2\alpha}(1-y)^{2}\over 1-x^{2\alpha}y}\biggr)+o(\theta^{2}),$ (3.5)
$\lambda_{1}\bigl(Y^{{1-\alpha\over 2}r}(XYX)^{\alpha r}Y^{{1-\alpha\over 2}r}\bigr)$ $=1+\theta^{2}\biggl(b_{11}+{b_{12}^{2}\over 1-x^{2\alpha r}y^{r}}\biggr)+o(\theta^{2}).$
Hence, for any $y>0$, whenever $x>0$ is sufficiently small, it must follows from (3.1) that
$r\biggl(-1+y+{x^{2\alpha}(1-y)^{2}\over 1-x^{2\alpha}y}\biggr)\leq b_{11}+{b_{12}^{2}\over 1-x^{2\alpha r}y^{r}}.$
As $x\searrow 0$, since
$a_{11}\to-\alpha r(1-y),\qquad a_{12}\to 0,$
$b_{11}\to-2\bigl(1-y^{{1-\alpha\over 2}r}\bigr)-\alpha r(1-y),\qquad b_{12}\to 1-y^{{1-\alpha\over 2}r},$
we must have
$-r+ry\leq-2\bigl(1-y^{{1-\alpha\over 2}r}\bigr)-\alpha r(1-y)+\bigl(1-y^{{1-\alpha\over 2}r}\bigr)^{2},\qquad y>0.$ (3.6)
Letting $y\searrow 0$ gives $r\leq-1-\alpha r$ so that $1\leq(1-\alpha)r$, i.e., $p/q\leq 1-\alpha$. Since $R_{\alpha,p}(A,B)=R_{1-\alpha,p}(B,A)$ and $SG_{\alpha,p}(A,B)=SG_{1-\alpha,p}(A,B)$ (see [41, Proposition 4.2(iii)]) for all $A,B>0$, note that (ii) implies condition (ii) with $1-\alpha$ in place of $\alpha$. Hence we have $p/q\leq\alpha$ too so that (iii) follows.
[…]
$\|SG_{\alpha,p}(A,B)\|_{\infty}\leq\|R_{\alpha,1}(A,B)\|_{\infty},$
or equivalently,
$B^{\alpha}\leq A^{\alpha-1}\implies SG_{\alpha,p}(A,B)\leq I.$ (3.7)
Now, we divide the proof into the two cases of $0<\alpha\leq 1/2$ and $1/2\leq\alpha<1$.
Case $0<\alpha\leq 1/2$. Since ${p\over\alpha}\leq 1$ by assumption, we have $B^{p}\leq A^{{\alpha-1\over\alpha}p}$ and hence
$A^{-p}\#B^{p}\leq A^{-p}\#A^{{\alpha-1\over\alpha}p}=A^{-{p\over 2\alpha}}.$
Since $2\alpha\leq 1$, we have $(A^{-p}\#B^{p})^{2\alpha}\leq A^{-p}$ so that $A^{p/2}(A^{-p}\#B^{p})^{2\alpha}A^{p/2}\leq I$. This gives $(A^{-p}\#B^{p})^{\alpha}A^{p}(A^{-p}\#B^{p})^{\alpha}\leq I$, showing (3.7).
Case $1/2\leq\alpha<1$. Since ${p\over 1-\alpha}\leq 1$, we have $B^{{\alpha\over 1-\alpha}p}\leq A^{-p}$ so that $A^{p}\leq B^{-{\alpha\over 1-\alpha}p}$. Hence
$B^{-p}\#A^{p}\leq B^{-p}\#B^{-{\alpha\over 1-\alpha}p}=B^{-{p\over 2(1-\alpha)}}.$
Since $2(1-\alpha)\leq 1$, we have $(B^{-p}\#A^{p})^{2(1-\alpha)}\leq B^{-p}$ so that $SG_{\alpha,p}(A,B)=SG_{1-\alpha,p}(B,A)\leq I$ as in the case $0<\alpha\leq 1/2$.
[…]
When $\alpha>1$, the next theorem says that we have a sufficient condition for $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ and that $R_{\alpha,q}\prec_{\log}SG_{\alpha,p}$ fails to hold for any $p,q>0$.
Theorem 3.4.
Let $\alpha>1$ and $p,q>0$.
  • (1)
    If $p/q\leq\alpha$, then $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ holds for all $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (2)
    For any $\alpha>1$ and any $p,q>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $R_{\alpha,q}(A,B)\not\prec_{\log}SG_{\alpha,p}(A,B)$.
Proof.
[…]
$(X^{\alpha}YX^{\alpha})^{r}\prec_{\log}Y^{{1-\alpha\over 2}r}(XYX)^{\alpha r}Y^{{1-\alpha\over 2}r}$
for all $X,Y>0$, where $r:=q/p$. To show this, it suffices to prove that
$(XYX)^{\alpha r}\leq Y^{(\alpha-1)r}\implies X^{\alpha}YX^{\alpha}\leq I.$ (3.8)
Now assume that $p/q\leq\alpha$ and so ${1\over\alpha r}\leq 1$. Then the inequality in the LHS of (3.8) implies that $XYX\leq Y^{\alpha-1\over\alpha}$ and hence
$X^{\alpha}YX^{\alpha}=X^{\alpha-1}(XYX)X^{\alpha-1}\leq X^{\alpha-1}Y^{\alpha-1\over\alpha}X^{\alpha-1}.$
Since $0<{\alpha-1\over\alpha}<1$, Araki’s log-majorization (Theorem 3.1(a)) gives
$X^{\alpha-1}Y^{\alpha-1\over\alpha}X^{\alpha-1}\prec_{\log}(X^{\alpha}YX^{\alpha})^{\alpha-1\over\alpha}.$
Therefore we have
$\|X^{\alpha}YX^{\alpha}\|_{\infty}\leq\big\|(X^{\alpha}YX^{\alpha})^{\alpha-1\over\alpha}\big\|_{\infty}=\|X^{\alpha}YX^{\alpha}\|_{\infty}^{\alpha-1\over\alpha},$
which implies that $\|X^{\alpha}YX^{\alpha}\|_{\infty}\leq 1$, showing (3.8).
[…]
$y^{(1-\alpha)r}-(1-\alpha)r(y-1)-1\leq 0,\qquad y>0.$ (3.9)
But this inequality is impossible since the above LHS goes to $\infty$ as $y\searrow 0$, so the result has been shown. ∎
Problem 3.5.
When $\alpha>1$, no necessary condition on $p,q>0$ under which $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ holds is known. In fact, when $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$, one has the reversed inequality of (3.9) but nothing follows from that. This suggests us that $SG_{\alpha,p}\prec_{\log}R_{\alpha,q}$ might hold for any $p,q>0$.
3.2 $\widetilde{S}G_{\alpha,p}\prec_{\log}R_{\alpha,q}$ and $R_{\alpha,q}\prec_{\log}\widetilde{S}G_{\alpha,p}$
The next theorem is concerned with the log-majorizations between $\widetilde{S}G_{\alpha,p}$ and $R_{\alpha,q}$ when $0<\alpha<1$.
Theorem 3.6.
Let $0<\alpha<1$ and $p,q>0$.
  • (1)
    The following conditions are equivalent:
    • (i)
      $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$;
    • (ii)
      $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$;
    • (iii)
      $p/q\leq\alpha$.
  • (2)
    If $\alpha\leq 1/2$ and $q\leq p$, then $R_{\alpha,q}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$ holds for all $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (3)
    Assume that $R_{\alpha,q}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$. Then we have $\alpha\leq 1/2$ and $p/q\geq 1/2$.
Proof.
[…]
$(X^{1/2}Y^{2(1-\alpha)}X^{1/2})^{r}\prec_{\log}Y^{{1-\alpha\over 2}r}(Y^{-1}\#_{1/\alpha}X)^{\alpha r}Y^{{1-\alpha\over 2}r},$
[…]
$\lambda_{1}^{r}(Y^{1-\alpha}XY^{1-\alpha})\leq\lambda_{1}\bigl(Y^{{1-\alpha\over 2}r}(Y^{-1}\#_{1/\alpha}X)^{\alpha r}Y^{{1-\alpha\over 2}r}\bigr)$ (3.10)
for all $X,Y\in\mathbb{M}_{2}^{++}$, noting that $\lambda(X^{1/2}Y^{2(1-\alpha)}X^{1/2})=\lambda(Y^{1-\alpha}XY^{1-\alpha})$.
Now let $Y:=A_{0}$ and $X:=B_{\theta}$ in (2.10) for $x,y>0$ with $xy\neq 1$, $x^{1-\alpha}y\neq 1$ and $x^{2(1-\alpha)}y\neq 1$. From the proof of [33, Theorem 4.35] we have
$Y^{1-\alpha}XY^{1-\alpha}=\begin{bmatrix}1-\theta^{2}(1-y)&\theta x^{1-\alpha}(1-y)\\ \theta x^{1-\alpha}(1-y)&x^{2(1-\alpha)}y+\theta^{2}x^{2(1-\alpha)}(1-y)\end{bmatrix}+o(\theta^{2})\quad\mbox{as $\theta\to 0$},$ (3.11)
and
$Y^{-1}\#_{1/\alpha}X=\begin{bmatrix}1+\theta^{2}u_{11}&\theta u_{12}\\ \theta u_{12}&x^{1-\alpha\over\alpha}y^{1\over\alpha}+\theta^{2}u_{22}\end{bmatrix}+o(\theta^{2}),$ (3.12)
where
$\begin{cases}u_{11}:=-{1-y\over\alpha}+{x(1-y)^{2}\bigl(1-\alpha-xy+\alpha x^{1\over\alpha}y^{1\over\alpha}\bigr)\over\alpha(1-xy)^{2}},\\ u_{12}:={(1-y)\bigl(1-x^{1\over\alpha}y^{1\over\alpha}\bigr)\over 1-xy}.\end{cases}$ (3.13)
Hence by [33, Example 3.2(1)] we compute
$(Y^{-1}\#_{1/\alpha}X)^{\alpha r}=\begin{bmatrix}1+\theta^{2}v_{11}&\theta v_{12}\\ \theta v_{12}&x^{(1-\alpha)r}y^{r}+\theta^{2}v_{22}\end{bmatrix}+o(\theta^{2}),$
where
$\begin{cases}v_{11}:=\alpha ru_{11}+{\alpha r-1-\alpha rx^{1-\alpha\over\alpha}y^{1\over\alpha}+x^{(1-\alpha)r}y^{r}\over\bigl(1-x^{1-\alpha\over\alpha}y^{1\over\alpha}\bigr)^{2}}\,u_{12}^{2},\\ v_{12}:={1-x^{(1-\alpha)r}y^{r}\over 1-x^{1-\alpha\over\alpha}y^{1\over\alpha}}\,u_{12},\end{cases}$
so that
$Y^{{1-\alpha\over 2}r}(Y^{-1}\#_{1/\alpha}X)^{\alpha r}Y^{{1-\alpha\over 2}r}=\begin{bmatrix}1+\theta^{2}v_{11}&\theta x^{{1-\alpha\over 2}r}v_{12}\\ \theta x^{{1-\alpha\over 2}r}v_{12}&x^{2(1-\alpha)r}y^{r}+\theta^{2}v_{22}\end{bmatrix}+o(\theta^{2})\quad\mbox{as $\theta\to 0$}.$ (3.14)
When $x^{1-\alpha}y<1$ and $x^{2(1-\alpha)y}<1$, applying [33, Lemma 3.6] to (3.11) and (3.14) we have
$\lambda_{1}^{r}(Y^{1-\alpha}XY^{1-\alpha})$ $=1+\theta^{2}r\biggl(-1+y+{x^{2(1-\alpha)}(1-y)^{2}\over 1-x^{2(1-\alpha)}y}\biggr)+o(\theta^{2}),$ (3.15)
$\lambda_{1}\bigl(Y^{{1-\alpha\over 2}r}(Y^{-1}\#_{1/\alpha}X)^{\alpha r}Y^{{1-\alpha\over 2}r}\bigr)$ $=1+\theta^{2}\biggl(v_{11}+{x^{(1-\alpha)r}v_{12}^{2}\over 1-x^{2(1-\alpha)r}y^{r}}\biggr)+o(\theta^{2}).$
Therefore, for any $x>0$, whenever $y>0$ is sufficiently small, it follows from (3.10) that
$r\biggl(-1+y+{x^{2(1-\alpha)}(1-y)^{2}\over 1-x^{2(1-\alpha)}y}\biggr)\leq v_{11}+{x^{(1-\alpha)r}v_{12}^{2}\over 1-x^{2(1-\alpha)r}y^{r}}.$
As $y\searrow 0$, since
$u_{11}\to-{1\over\alpha}+{1-\alpha\over\alpha}\,x,\qquad u_{12}\to 1,$
$v_{11}\to-r+r(1-\alpha)x+\alpha r-1,\qquad v_{12}\to 1,$
we must have
$rx^{2(1-\alpha)}\leq r(1-\alpha)x+\alpha r-1+x^{(1-\alpha)r},\qquad x>0.$ (3.16)
Letting $x\searrow 0$ gives $0\leq\alpha r-1$, i.e., $1/r\leq\alpha$. Hence (iii) follows.
[…]
$B^{\alpha q}\leq A^{(\alpha-1)q}\implies\widetilde{S}G_{\alpha,p}(A,B)\leq I.$ (3.17)
Assume that $B^{\alpha q}\leq A^{(\alpha-1)q}$; then we have $B^{p}\leq A^{{\alpha-1\over\alpha}p}$ since ${p\over\alpha q}\leq 1$. Therefore,
$A^{-p}\#_{\alpha}B^{p}\leq A^{-p}\#_{\alpha}A^{{\alpha-1\over\alpha}p}=A^{-2(1-\alpha)p},$
so that we have $A^{2(1-\alpha)p}\leq(A^{-p}\#_{\alpha}B^{p})^{-1}$ and $(A^{-p}\#_{\alpha}B^{p})^{1/2}A^{2(1-\alpha)p}(A^{-p}\#_{\alpha}B^{p})^{1/2}\leq I$, showing (3.17).
(2) When $\alpha=1/2$, since $\lambda(\widetilde{S}G_{1/2,p}(A,B))=\lambda(SG_{1/2,p}(A,B))=\lambda(R_{1/2,2p}(A,B))$, the assertion in this case follows from Theorem 3.1(a) if $q\leq 2p$ (weaker than $q\leq p$). Next assume that $\alpha<1/2$ and $q\leq p$. By Theorem 3.1(a) again we may show that $R_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$, which is more explicitly written as
$A^{{1-\alpha\over 2}p}B^{\alpha p}A^{{1-\alpha\over 2}p}\prec_{\log}A^{{1-2\alpha\over 2}p}(A^{p/2}B^{p}A^{p/2})^{\alpha}A^{{1-2\alpha\over 2}p}.$ (3.18)
Set $p_{1}:={1\over 1-2\alpha}$, $q_{1}:={\alpha\over 1-2\alpha}=\alpha p_{1}$, $A_{1}:=A^{(1-2\alpha)p}$ and $B_{1}:=B^{(1-2\alpha)p}$. Since
$A^{p}=A_{1}^{p_{1}},\qquad B^{p}=B_{1}^{p_{1}},\qquad A^{(1-\alpha)p}=A_{1}^{1-\alpha\over 1-2\alpha}=A_{1}^{1+q_{1}},\qquad B^{\alpha p}=B_{1}^{\alpha\over 1-2\alpha}=B_{1}^{q_{1}},$
the log-majorization in (3.18) is equivalently written as
$A_{1}^{1+q_{1}\over 2}B_{1}^{q_{1}}A_{1}^{1+q_{1}\over 2}\prec_{\log}A_{1}^{1/2}\bigl(A_{1}^{p_{1}/2}B_{1}^{p_{1}}A_{1}^{p_{1}/2}\bigr)^{q_{1}/p_{1}}A_{1}^{1/2},$
which is exactly the BLP log-majorization due to Bebiano–Lemos–Providência [6, Theorem 2.1].
[…]
$rx^{2(1-\alpha)}\geq r(1-\alpha)x+\alpha r-1+x^{(1-\alpha)r},\qquad x>0,$
where $r:=q/p$. Letting $x\searrow 0$ gives $0\geq\alpha r-1$, i.e., $1/r\geq\alpha$. Also, looking at the order as $x\to\infty$ we have $2(1-\alpha)\geq 1$ and $2(1-\alpha)\geq(1-\alpha)r$, so that $\alpha\leq 1/2$ and $1/r\geq 1/2$. Hence the result follows. ∎
The next theorem is concerned with the log-majorizations between $\widetilde{S}G_{\alpha,p}$ and $R_{\alpha,q}$ when $\alpha>1$.
Theorem 3.7.
Let $\alpha>1$ and $p,q>0$.
  • (1)
    For any $p,q>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\widetilde{S}G_{\alpha,p}(A,B)\not\prec_{\log}R_{\alpha,q}(A,B)$.
  • (2)
    If $p/q\geq\alpha$, then $R_{\alpha,q}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$ holds for all $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (3)
    Assume that $R_{\alpha,q}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$. Then we have $p/q\geq 1/2$.
Proof.
[…]
$\widetilde{S}G_{\alpha,p}(A,B)\leq I\implies B^{\alpha q}\leq A^{(\alpha-1)q}.$
If $\widetilde{S}G_{\alpha,p}(A,B)\leq I$, then one has $A^{-p}\#_{\alpha}B^{p}\leq A^{2(\alpha-1)p}$, so that $(A^{p/2}B^{p}A^{p/2})^{\alpha}\leq A^{(2\alpha-1)p}$. Since $\alpha>1$, this gives $A^{p/2}B^{p}A^{p/2}\leq A^{{2\alpha-1\over\alpha}p}$ and hence $B^{p}\leq A^{{\alpha-1\over\alpha}p}$. Since ${\alpha q\over p}\leq 1$, we have $B^{\alpha q}\leq A^{(\alpha-1)q}$.
[…]
$rx^{2(1-\alpha)}\geq r(1-\alpha)x+\alpha r-1+x^{(1-\alpha)r},\qquad x>0,$
where $r:=q/p$. Now assume that $p/q<1/2$. Then, since $0>2(1-\alpha)>(1-\alpha)r$, the above inequality is impossible in the limit $x\searrow 0$. Hence $p/q\geq 1/2$ must hold. ∎
Problem 3.8.
When $0<\alpha<1$, there is a gap between the sufficient condition in Theorem 3.6(2) and the necessary condition in Theorem 3.6(3) for $R_{\alpha,q}\prec_{\log}\widetilde{S}G_{\alpha,p}$. When $\alpha>1$, there is also a big gap between the sufficient condition in Theorem 3.7(2) and the necessary condition in Theorem 3.7(3) for $R_{\alpha,q}\prec_{\log}\widetilde{S}G_{\alpha,p}$. Hence the problem of characterizing $R_{\alpha,q}\prec_{\log}\widetilde{S}G_{\alpha,p}$ is far from completed. When $\alpha=1/2$, note that the condition in Theorem 3.6(2) is not sharp as mentioned in its proof, while the sufficient condition in Theorem 3.7(2) is sharp.
Remark 3.9.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p,q>0$. For any pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ as mentioned in Remark 2.4, note that $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$ if and only if $\lambda_{1}(\mathcal{M}_{\alpha,p}(A,B))\leq\lambda_{1}(\mathcal{N}_{\alpha,q}(A,B))$ for all $A,B\in\mathbb{M}_{2}^{++}$. Therefore, one can easily find a necessary condition for $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ by use of expressions given in Lemma 2.5. For instance, if $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$, then from (2.11) and (2.13) one must have
${2\alpha(1-\alpha)\over p}\cdot{1-x^{p}\over 1+x^{p}}\leq{-1-x^{q}+x^{\alpha q}+x^{(1-\alpha)q}\over q(1-x^{q})},\qquad x\in(0,1).$
Letting $x\searrow 0$ gives $p/q\leq 2\alpha(1-\alpha)$ if $0<\alpha<1$. Similarly, if $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$, then one has $p/q\geq 2\alpha(1-\alpha)$ if $0<\alpha<1$, and a contradiction if $\alpha>1$. Hence Theorem 3.4(2) follows again, while we have necessary conditions weaker than Theorem 3.3 since $\min\{\alpha,1-\alpha\}\leq 2\alpha(1-\alpha)\leq\max\{\alpha,1-\alpha\}$ for $0<\alpha<1$.
Also, if $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$, then from (2.11) and (2.14) one has
${1\over p}\biggl\{\alpha\,{1-x^{p}\over 1+x^{p}}+{x^{p}+x^{2p}-x^{(2\alpha+1)p}-x^{2(1-\alpha)p}\over(1-x^{p})(1+x^{p})^{2}}\biggr\}\leq{-1-x^{q}+x^{\alpha q}+x^{(1-\alpha)q}\over q(1-x^{q})},\qquad x\in(0,1).$
[…]
3.3 Regarding $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$, $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$
While the log-majorizations $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$ for $\alpha\in(0,2]\setminus\{1\}$ were completely characterized as stated in Theorem 3.1(b) and (c), the characterizations of $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$, as well as $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$ for $\alpha>2$, have not been obtained yet. In this subsection we discuss the problem though without complete success.
The proof of Theorem 3.1(d) for $\alpha\geq 2$ was given in [31] in an indirect way combining (e) and (f), so the sufficient condition in (d) is probably not best possible. In a similar way, a certain sufficient condition for $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ (resp., $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$) is given as follows by combining the log-majorizations in Section 3.1 (resp., Section 3.2).
Proposition 3.10.
Let $0<\alpha<1$ and $p,q>0$.
  • (1)
    If $p/q\leq\min\{{\alpha\over 1-\alpha},{1-\alpha\over\alpha}\}$, then we have $SG_{\alpha,p}(A,B)\prec_{\log}SG_{\alpha,q}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (2)
    If $\alpha\leq 1/2$ and $p/q\leq\alpha$, then we have $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$.
Proof.
(1) Assume that $0<\alpha\leq 1/2$ and $p/q\leq{\alpha\over 1-\alpha}$. Since ${q\over p/\alpha}\geq 1-\alpha$, by Theorem 3.3 we have
$SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,{p/\alpha}}(A,B)\prec_{\log}SG_{\alpha,q}(A,B)$
Assume that $1/2\leq\alpha<1$ and $p/q\leq{1-\alpha\over\alpha}$. Since ${q\over p/(1-\alpha)}\geq\alpha$, we have similarly
$SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,{p/(1-\alpha)}}(A,B)\prec_{\log}SG_{\alpha,q}(A,B).$
(2) Assume that $\alpha\leq 1/2$ and $p/\alpha\leq q$. Then by Theorem 3.6 we have
$\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,p/\alpha}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B).$
The next proposition shows that $p\leq q$ or $p\geq q$ is anyhow necessary for the log-majorizations of our concern to hold.
Proposition 3.11.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p,q>0$.
  • (1)
    Assume that $G_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$. Then we have $p\geq q$ if $0<\alpha<1$, and $p\leq q$ if $\alpha>1$.
  • (2)
    Assume that $SG_{\alpha,p}(A,B)\prec_{\log}SG_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$. Then we have $p\leq q$ if $0<\alpha<1$, and $p\geq q$ if $\alpha>1$.
  • (3)
    Assume that $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$ holds for all $A,B\in\mathbb{M}_{2}^{++}$. Then we have $p\leq q$ if $0<\alpha<1$, and $p\geq q$ if $\alpha>1$.
Proof.
These are easily checked by use of expressions (2.12)–(2.14) in Lemma 2.5 similarly to the discussions in Remark 3.9. The details are left to the reader. ∎
Problem 3.12.
What we are most interested in is whether a necessary condition $p\leq q$ or $p\geq q$ confirmed in Proposition 3.11 is indeed sufficient for $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$ to hold or not. The problem is to show the variants of Ando–Hiai’s inequality [3] for $SG_{\alpha,p}$ and $\widetilde{S}G_{\alpha,q}$. Note that the condition on $p,q$ in Proposition 3.11(1) and that in Proposition 3.11(2) and (3) are opposite. This suggests us that two types of weighted spectral geometric means are quite different from Kubo–Ando’s weighted geometric mean $_{\alpha}$.
3.4 More log-majorizations
In this subsection we will examine log-majorizations between $G_{\alpha}$, $SG_{\alpha}$ and $\widetilde{S}G_{\alpha}$ in an easy way of putting $R_{\alpha}$ in the middle of two of them for sufficient conditions, or the outside of the two for necessary conditions, otherwise by use of Lemma 2.5 as in Remark 3.9.
Proposition 3.13.
  • (1)
    Let $0<\alpha<1$. For any $p,q>0$ we have $G_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (2)
    Let $\alpha>1$. For any $p,q>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $G_{\alpha,q}(A,B)\not\prec_{\log}SG_{\alpha,p}(A,B)$.
  • (3)
    Let $\alpha>1$. If $p/q\leq\max\bigl\{2,{\alpha\over\alpha-1}\bigr\}$, then we have $SG_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,q}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$.
Proof.
(1) Thanks to Theorem 3.2(1) and 3.3(2), taking an $r>0$ with $p/r\geq\max\{\alpha,1-\alpha\}$ we have
$G_{\alpha,q}(A,B)\prec_{\log}R_{\alpha,r}(A,B)\prec_{\log}SG_{\alpha,p}(A,B).$
(2) is easily verified by using expressions (2.12) and (2.13) in Lemma 2.5 similarly to the discussions in Remark 3.9.
(3) Assume that $\alpha>1$ and $p/q\leq\min\bigl\{2,{\alpha\over\alpha-1}\bigr\}$. For every $A,B\geq 0$ with $s(A)\geq s(B)$, letting $r:=p/\alpha$, we have $SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,r}(A,B)$ by Theorem 3.4(1). Moreover, since $q/r=q\alpha/p\geq\max\{\alpha/2,\alpha-1\}$, we have $R_{\alpha,r}(A,B)\prec_{\log}G_{\alpha,q}(A,B)$ by Theorem 3.2(2). Therefore, $SG_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,q}(A,B)$. ∎
Proposition 3.14.
  • (1)
    For any $\alpha\in(0,\infty)\setminus\{1\}$ and any $p,q>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\widetilde{S}G_{\alpha,p}(A,B)\not\prec_{\log}G_{\alpha,q}(A,B)$.
  • (2)
    Let $0<\alpha\leq 1/2$. For any $p,q$ we have $G_{\alpha,q}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (3)
    Let $\alpha>1$. If $q/p\leq\min\bigl\{1/2,{\alpha-1\over\alpha}\bigr\}$, then we have $G_{\alpha,q}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$.
Proof.
(1) Assume that $\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$. When $0<\alpha<1$, by Theorem 3.2(1) we have
$\widetilde{S}G_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,q}(A,B)\prec_{\log}R_{\alpha,r}(A,B)$
for any $r>0$ and all $A,B\in\mathbb{M}_{2}^{++}$, contradicting Theorem 3.6(1). (The $0<\alpha<1$ case is seen also by use of expressions (2.12) and (2.14).) When $\alpha>1$, by Theorem 3.2(1) again the above holds for any sufficiently large $r>0$ and all $A,B\in\mathbb{M}_{2}^{++}$, contradicting Theorem 3.7(1).
(2) Let $0<\alpha\leq 1/2$. For any $p,q>0$, by Theorems 3.2(1) and 3.6(2) we have
$G_{\alpha,q}(A,B)\prec_{\log}R_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$
for all $A,B\geq 0$ with $s(A)\geq s(B)$.
(3) Assume that $\alpha>1$ and $q/p\leq\min\bigl\{1/2,{\alpha-1\over\alpha}\bigr\}$. Since $q/(p/\alpha)\leq\min\{\alpha/2,\alpha-1\}$, by Theorems 3.2(1) and 3.7(2) we have
$G_{\alpha,q}(A,B)\prec_{\log}R_{\alpha,p/\alpha}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$
for all $A,B\geq 0$ with $s(A)\geq s(B)$. ∎
Proposition 3.15.
  • (1)
    Let $0<\alpha<1$ and $p,q>0$. If $SG_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$, then we have $\alpha\leq 1/2$ and $p/q\leq 2(1-\alpha)$.
  • (2)
    Let $0<\alpha<1$. If $p/q\geq\max\bigl\{1,{1-\alpha\over\alpha}\bigr\}$, then $\widetilde{S}G_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$. Morevoer, if $\widetilde{S}G_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$, then we have $p/q\geq 2(1-\alpha)$. (Note that $2(1-\alpha)\leq\max\bigl\{1,{1-\alpha\over\alpha}\bigr\}$ for $0<\alpha<1$.)
  • (3)
    Let $\alpha>1$. If $p\leq q$, then $SG_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (4)
    Let $\alpha>1$. For any $p,q>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\widetilde{S}G_{\alpha,q}(A,B)\not\prec_{\log}SG_{\alpha,p}(A,B)$.
Proof.
(1) Assume that $SG_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$. For $r:=p/\max\{\alpha,1-\alpha\}$, by Theorem 3.3(2) we have
$R_{\alpha,r}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$
for all $A,B\in\mathbb{M}_{2}^{++}$. Hence by Theorem 3.6(3) we must have $\alpha\leq 1/2$ and $q/r\geq 1/2$ so that $p/q\leq 2\max\{\alpha,1-\alpha\}=2(1-\alpha)$.
(2) Assume that $0<\alpha<1$ and $p/q\geq\max\bigl\{1,{1-\alpha\over\alpha}\bigr\}$. Let $r:=q/\alpha$ and so $p/r\geq\max\{\alpha,1-\alpha\}$. Thanks to Theorem 3.6(1) and 3.3(2) we have
$\widetilde{S}G_{\alpha,q}(A,B)\prec_{\log}R_{\alpha,r}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$
for all $A,B\geq 0$ with $s(A)\geq s(B)$. Next, assume that $\widetilde{S}G_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ for all $A,B\in\mathbb{M}_{2}^{++}$. Then $p/q\geq 2(1-\alpha)$ is easily verified from expressions (2.13) and (2.14).
(3) Letting $r:=p/\alpha$ and so $q/r\geq\alpha$, by Theorems 3.4(1) and 3.7(2) we have
$SG_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,r}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,q}(A,B)$
for all $A,B\geq 0$ with $s(A)\geq s(B)$.
(4) is easily verified from expressions (2.13) and (2.14) again. ∎
The section ends with log-majorizations of $LE_{\alpha}$ with other quasi-geometric type means. The next two remarks are easy but convenient.
Remark 3.16.
Let $\mathcal{M}\in\{R,G,SG,\widetilde{S}G\}$, $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$.
(1) The log-majorization $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}LE_{\alpha}(A,B)$ (resp., $LE_{\alpha}(A,B)\prec_{\log}\mathcal{M}_{\alpha,p}(A,B)$) holds for all $A,B>0$ if and only if $\mathcal{M}_{\alpha,1}(A,B)\prec_{\log}LE_{\alpha}(A,B)$ (resp., $LE_{\alpha}(A,B)\prec_{\log}\mathcal{M}_{\alpha,1}(A,B)$) for all $A,B>0$. Indeed, if $\mathcal{M}_{\alpha,1}(A,B)\prec_{\log}LE_{\alpha}(A,B)$ for all $A,B>0$, then
$\mathcal{M}_{\alpha,1}(A^{p},B^{p})\prec_{\log}LE_{\alpha}(A^{p},B^{p})=LE_{\alpha}(A,B)^{p}$
so that we have $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}LE_{\alpha}(A,B)$ for all $A,B>0$.
(2) If $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}LE_{\alpha}(A,B)$ (resp., $LE_{\alpha}(A,B)\prec_{\log}\mathcal{M}_{\alpha,p}(A,B)$) holds for all $A,B\in\mathbb{M}_{2}^{++}$, then the reverse log-majorization $LE_{\alpha}(A,B)\prec_{\log}\mathcal{M}_{\alpha,p}(A,B)$ (resp., $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}LE_{\alpha}(A,B)$) does not hold for some $A,B\in\mathbb{M}_{2}$. Otherwise, we must have $\lambda(\mathcal{M}_{\alpha,p}(A,B))=\lambda(LE_{\alpha}(A,B))$ for all $A,B\in\mathbb{M}_{2}^{++}$, which contradicts the assertion of Remark 2.4(4).
The known cases are summarized in the following proposition.
Proposition 3.17.
  • (a)
    $LE_{\alpha}(A,B)\prec_{\log}R_{\alpha,p}(A,B)$ for all $\alpha\in(0,\infty)\setminus\{1\}$, $p>0$ and $A,B\geq 0$ (with $s(A)\geq s(B)$ for $\alpha>1$).
  • (b)
    $G_{\alpha,p}(A,B)\prec_{\log}LE_{\alpha}(A,B)$ for all $\alpha\in(0,1)$, $p>0$ and $A,B\geq 0$.
  • (c)
    $LE_{\alpha}(A,B)\prec_{\log}G_{\alpha,p}(A,B)$ for all $\alpha\in(1,2]$, $p>0$ and $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (d)
    $LE_{\alpha}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ for all $\alpha\in(0,1)$, $p>0$ and $A,B\geq 0$ with $s(A)\geq s(B)$.
See [3, 11] for (a) and (b), [39] for (c), and [23, 22] for (d). When $0<\alpha<1$, it is worth noting that for every (a) and (b) above are supplemented with the quasi-arithmetic and the quasi-harmonic means as follows: if $r\geq p/2$ and $s\geq q$ then
$\mathcal{H}_{\alpha,s}(A,B)\leq_{\lambda}G_{\alpha,q}(A,B)\prec_{\log}LE_{\alpha}(A,B)\prec_{\log}R_{\alpha,p}(A,B)\leq_{\lambda}\mathcal{A}_{\alpha,r}(A,B),\qquad A,B\geq 0,$
where the additional inequalities in both sides are seen from [33, Propositions 4.9, 4.19 and Remark 2.7(3)].
From log-majorizations given in Sections 3.1 and 3.2 we have more results as follows:
Corollary 3.18.
  • (1)
    $LE_{\alpha}(A,B)\prec_{\log}G_{\alpha,p}(A,B)$ for all $\alpha>1$, $p>0$ and $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (2)
    For any $\alpha>1$ and $p>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $LE_{\alpha}(A,B)\not\prec_{\log}SG_{\alpha,p}(A,B)$.
  • (3)
    For any $0<\alpha<1$ and $p>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\widetilde{S}G_{\alpha,p}(A,B)\not\prec_{\log}LE_{\alpha}(A,B)$.
  • (4)
    $LE_{\alpha}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$ for all $\alpha\in(0,1/2]$, $p>0$ and $A,B\geq 0$ with $s(A)\geq s(B)$.
  • (5)
    $LE_{\alpha}(A,B)\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$ for all $\alpha>1$, $p>0$ and $A,B\geq 0$ with $s(A)\geq s(B)$.
Proof.
(1) follows by letting $p\searrow 0$ in (c) and (d) of Theorem 3.1. (2) and (3) are immediately seen by comparing (2.15) with (2.13) and (2.14), respectively. (4) and (5) follow as the $q\searrow 0$ limits (due to Theorem 2.2) of the log-majorizations in Theorems 3.6(2) and 3.7(2), respectively. ∎
The next theorem says that $LE_{\alpha}\prec_{\log}\widetilde{S}G_{\alpha,p}$ for $1/2<\alpha<1$ fails to hold.
Theorem 3.19.
For any $\alpha\in(1/2,1)$ and $p>0$ there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $LE_{\alpha}(A,B)\not\prec_{\log}\widetilde{S}G_{\alpha,p}(A,B)$.
Proof.
[…]
$LE_{\alpha}(Y,Y^{-1}\#_{1/\alpha}X)\prec_{\log}X^{1/2}Y^{2(1-\alpha)}X^{1/2},\qquad X,Y\in\mathbb{M}_{2}^{++},$
equivalently,
$\exp\lambda_{1}((1-\alpha)\log Y+\alpha\log(Y^{-1}\#_{1/\alpha}X))\leq\lambda_{1}(Y^{1-\alpha}XY^{1-\alpha}),\qquad X,Y\in\mathbb{M}_{2}^{++}.$ (3.19)
Now let $Y:=A_{0}$ and $X:=B_{\theta}$ in (2.10) for $x,y>0$ with $xy<1$ and $x^{2(1-\alpha)}y<1$. By (3.15) (with $r=1$) we have
$\lambda_{1}(Y^{1-\alpha}XY^{1-\alpha})=1+\theta^{2}\biggl(-1+y+{x^{2(1-\alpha)}(1-y)^{2}\over 1-x^{2(1-\alpha)}y}\biggr)+o(\theta^{2})\quad\mbox{as $\theta\to 0$}.$ (3.20)
On the other hand, apply [33, Example 3.2(3)] to (3.12) and (3.13) to compute
$\log(Y^{-1}\#_{1/\alpha}X)=\begin{bmatrix}\theta^{2}w_{11}&\theta w_{12}\\ \theta w_{12}&\log x^{1-\alpha\over\alpha}y^{1\over\alpha}+\theta^{2}w_{22}\end{bmatrix}+o(\theta^{2}),$
where
$\begin{cases}w_{11}:=u_{11}+{1-x^{1-\alpha\over\alpha}y^{1\over\alpha}+\log x^{1-\alpha\over\alpha}y^{1\over\alpha}\over\bigl(1-x^{1-\alpha\over\alpha}y^{1\over\alpha}\bigr)^{2}}\,u_{12}^{2},\\ w_{12}:=-{\log x^{1-\alpha\over\alpha}y^{1\over\alpha}\over 1-x^{1-\alpha\over\alpha}y^{1\over\alpha}}\,u_{12}.\end{cases}$
Hence it follows that
$(1-\alpha)\log Y+\alpha\log(Y^{-1}\#_{1/\alpha}X)=\begin{bmatrix}\theta^{2}\alpha w_{11}&\theta\alpha w_{12}\\ \theta\alpha w_{12}&\log x^{2(1-\alpha)}y+\theta^{2}\alpha w_{22}\end{bmatrix}+o(\theta^{2}),$
so that by [33, Lemma 3.6],
$\lambda_{1}((1-\alpha)\log Y+\alpha\log(Y^{-1}\#_{1/\alpha}X))=\theta^{2}\biggl(\alpha w_{11}-{\alpha^{2}w_{12}^{2}\over\log x^{2(1-\alpha)}y}\biggr)+o(\theta^{2}),$
which implies that
$\exp\lambda_{1}((1-\alpha)\log Y+\alpha\log(Y^{-1}\#_{1/\alpha}X))=1+\theta^{2}\biggl(\alpha w_{11}-{\alpha^{2}w_{12}^{2}\over\log x^{2(1-\alpha)}y}\biggr)+o(\theta^{2}).$ (3.21)
Therefore, by (3.19)–(3.21) we must have
$\alpha w_{11}-{\alpha^{2}w_{12}^{2}\over\log x^{2(1-\alpha)}y}\leq-1+y+{x^{2(1-\alpha)}(1-y)^{2}\over 1-x^{2(1-\alpha)}y}$ (3.22)
for any $x>0$, whenever $y>0$ is sufficiently small. A direct computation yields that the LHS of (3.22) is arranged as
$\alpha u_{11}+{2\alpha\big(1-x^{1-\alpha\over\alpha}y^{1\over\alpha}\bigr)\log x^{1-\alpha}+\log^{2}x^{1-\alpha}-\alpha x^{1-\alpha\over\alpha}y^{1\over\alpha}\log y+(\alpha+\log x^{1-\alpha})\log y\over\bigl(1-x^{1-\alpha\over\alpha}y^{1\over\alpha}\bigr)^{2}(2\log x^{1-\alpha}+\log y)}\,u_{12}^{2},$
whose limit as $y\searrow 0$ is equal to $1+(1-\alpha)x+\alpha+\log x^{1-\alpha}$, since $u_{11}\to-{1\over\alpha}+{1-\alpha\over\alpha}\,x$ and $u_{12}\to 1$ as $y\searrow 0$ thanks to (3.13). Hence, letting $y\searrow 0$ in (3.22) gives
$(1-\alpha)x+\alpha+\log x^{1-\alpha}\leq x^{2(1-\alpha)},\qquad x>0,$
which implies that $1\leq 2(1-\alpha)$, i.e., $\alpha\leq 1/2$, as desired. ∎
Problem 3.20.
The case remaining open is whether $SG_{\alpha,p}\prec_{\log}LE_{\alpha}$ for $\alpha>1$. Note that this is affirmative if $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ if $\alpha>1$ and $p\geq q$ (see Problem 3.12).
For the convenience of the reader, Proposition 3.17, Corollary 3.18 (also Remark 3.16) and Theorem 3.19 together are summarized as follows. Here, “all pp” says that the designated log-majorization holds for all $A,B>0$ and all $p>0$, and “none” says that for any $p>0$ it fails for some $A,B\in\mathbb{M}_{2}^{++}$. An asymmetric behavior of $\widetilde{S}G_{\alpha,p}$ for $0<\alpha<1$ around $\alpha=1/2$ is worth noting, that is a reflection of asymmetry of $\widetilde{S}G_{\alpha,p}$ under interchanging $\alpha$ and $1-\alpha$.
$0<\alpha<1$ $\alpha>1$///
$R_{\alpha,p}\prec_{\log}LE_{\alpha}$ none none
$LE_{\alpha}\prec_{\log}R_{\alpha,p}$ all $p$ all $p$
$G_{\alpha,p}\prec_{\log}LE_{\alpha}$ all $p$ none
$LE_{\alpha}\prec_{\log}G_{\alpha,p}$ none all $p$
$SG_{\alpha,p}\prec_{\log}LE_{\alpha}$ none    ?
$LE_{\alpha}\prec_{\log}SG_{\alpha,p}$ all $p$ none
$\widetilde{S}G_{\alpha,p}\prec_{\log}LE_{\alpha}$ none none
$LE_{\alpha}\prec_{\log}\widetilde{S}G_{\alpha,p}$ $\thinspace\begin{array}[]{ll}\mbox{all $p$\, for $0<\alpha\leq 1/2$}\\ \mbox{none for $1/2<\alpha<1$}\end{array}$ all $p$
4 Equality cases in norm inequalities derived from log-majorizations
Recall that if $X,Y\in\mathbb{M}_{n}^{+}$ and $X\prec_{\log}Y$, then $\|X\|\leq\|Y\|$ holds for all unitarily invariant norms $\|\cdot\|$ on $\mathbb{M}_{n}$; see, e.g., [29, Proposition 4.4.13]. Therefore, if a pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,p})$ of quasi-geometric type means satisfies $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$, then we have $\|\mathcal{M}_{\alpha,p}(A,B)\|\leq\|\mathcal{N}_{\alpha,q}(A,B)\|$ for all $A,B\geq 0$ and for any unitarily invariant norm $\|\cdot\|$. In this section we aim at characterizing the equality case in the above norm inequality in terms of commutativity $AB=BA$ in certain cases of the pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,p})$ and the norm $\|\cdot\|$. This is reasonable since $AB=BA$ implies that $\mathcal{M}_{\alpha,p}(A,B)=\mathcal{N}_{\alpha,q}(A,B)$ holds for any pair of quasi-geometric type means. We say that a norm $\|\cdot\|$ on $\mathbb{M}_{n}$ is strictly increasing if $A\geq B\geq 0$ and $\|A\|=\|B\|$ imply $A=B$. The Schatten $p$-norms where $1\leq p<\infty$ (particularly, the trace norm) are typical examples of strictly increasing unitarily invariant norms.
Theorem 4.1.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p,q>0$. Consider the following pairs of quasi-geometric matrix means:
$(R_{\alpha,p},R_{\alpha,q})\quad\mbox{for}\ \alpha\in(0,\infty)\setminus\{1\},\ p\neq q,$ (4.1)
$(G_{\alpha,p},R_{\alpha,q})\quad\mbox{for}\ \begin{cases}0<\alpha<1,\ p,q>0,\ \mbox{or}\\ \alpha>1,\ p/q<\min\{\alpha/2,\alpha-1\},\ \mbox{or}\\ \alpha>1,\ p/q>\max\{\alpha/2,\alpha-1\},\end{cases}$ (4.2)
$(SG_{\alpha,p},R_{\alpha,q})\quad\mbox{for}\ \begin{cases}0<\alpha<1,\ p/q>\max\{\alpha,1-\alpha\},\ \mbox{or}\\ 0<\alpha<1,\ p/q<\min\{\alpha,1-\alpha\},\ \mbox{or}\\ \alpha>1,\ p/q<\alpha,\end{cases}$ (4.3)
$(\widetilde{S}G_{\alpha,p},R_{\alpha,q})\quad\mbox{for}\ \begin{cases}0<\alpha<1,\ p/q<\alpha,\ \mbox{or}\\ 0<\alpha\leq 1/2,\ q<p,\ \mbox{or}\\ \alpha>1,\ p/q>\alpha,\end{cases}$ (4.4)
$(LE_{\alpha},R_{\alpha,p})\quad\mbox{for}\ \alpha\in(0,\infty)\setminus\{1\},\ p>0,$ (4.5)
$(G_{\alpha,p},G_{\alpha,q})\quad\mbox{for}\ \alpha\in(0,2]\setminus\{1\},\ p\neq q,$ (4.6)
$(SG_{\alpha,p},G_{\alpha,q})\quad\mbox{for}\ \begin{cases}0<\alpha<1,\ p,q>0,\ \mbox{or}\\ 1<\alpha\leq 2,\ p/q<\max\bigl\{2,{\alpha\over\alpha-1}\bigr\},\end{cases}$ (4.7)
$(\widetilde{S}G_{\alpha,p},G_{\alpha,q})\quad\mbox{for}\ \begin{cases}0<\alpha\leq 1/2,\ p,q>0,\ \mbox{or}\\ 1<\alpha\leq 2,\ q/p<\min\bigl\{1/2,{\alpha-1\over\alpha}\bigr\},\end{cases}$ (4.8)
$(LE_{\alpha},G_{\alpha,p})\quad\mbox{for}\ \alpha\in(0,2]\setminus\{1\},\ p>0.$ (4.9)
Let $\|\cdot\|$ be any strictly increasing unitarily invariant norm. Then the following hold:
  • (1)
    Let $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ be any of (4.1)–(4.5) and $A,B\geq 0$, with an additional assumption $s(A)\geq s(B)$ for the $\alpha>1$ case of (4.1) and (4.2) and for (4.3) and (4.3). If $\|\mathcal{M}_{\alpha,p}(A,B)\|=\|\mathcal{N}_{\alpha,q}(A,B)\|$, then $AB=BA$
  • (2)
    Let $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ be any of (4.6)–(4.9) and $A,B>0$. If $\|\mathcal{M}_{\alpha,p}(A,B)\|=\|\mathcal{N}_{\alpha,q}(A,B)\|$, then $AB=BA$
Proof.
(1) The main of this part is the assertion for (4.1), which was formerly shown in [28, Theorem 2.1] and the proof has been updated in [32, Appendix A]. [^2] We note that the claim [32, (A.3)] and its reasoning in the last five lines of [32, p. 544] is incorrect. But (A.3) is unnecessary in the proof of [32, Appendix]. Therefore, these lines and Remark A.2 of [32] should be deleted. The other cases in (4.2)–(4.5) are all reduced to (4.1). Indeed, assume that $\|G_{\alpha,p}(A,B)\|=\|R_{\alpha,q}(A,B)\|$ for $\alpha,p,q$ satisfying one of the three conditions in (4.2). When either $0<\alpha<1$ with $p,q>0$ arbitrary, or $\alpha>1$ and $p/q<\min\{\alpha/2,\alpha-1\}$, choose a $q^{\prime}\in(0,q)$ such that $\alpha,p,q^{\prime}$ satisfy the same condition. By Theorem 3.1(a) and (e) one has
$\|G_{\alpha,p}(A,B)\|\leq\|R_{\alpha,q^{\prime}}(A,B)\|\leq\|R_{\alpha,q}(A,B)\|.$
Hence $\|R_{\alpha,q^{\prime}}(A,B)\|=\|R_{\alpha,q}(A,B)\|$, which gives $AB=BA$ by the (4.1) case. When $\alpha>1$ and $p/q>\max\{\alpha/2,\alpha-1\}$, choose a $q^{\prime}>q$ such that $\alpha,p,q^{\prime}$ satisfy the same condition. By Theorem 3.1(a) and (f) one has
$\|R_{\alpha,q}(A,B)\|\leq\|R_{\alpha,q^{\prime}}(A,B)\|\leq\|G_{\alpha,p}(A,B)\|,$
so that $\|R_{\alpha,q}(A,B)\|=\|R_{\alpha,q^{\prime}}(A,B)\|$ and $AB=BA$ follows. The proof is similar for (4.3) by using Theorems 3.1(g) (or 3.3(2)), 3.3(1) and 3.4(1) for the three cases in (4.3), respectively. The proof for (4.4) is also similar by Theorems 3.6(1), 3.6(2) and 3.7(2) for the three cases in (4.4), respectively. For (4.5) we may use Proposition 3.17(a).
(2) The main of this part is the assertion for (4.6), which was formerly shown for $0<\alpha<1$ in [28, Theorem 3.1]. The assertion for $1<\alpha\leq 2$ is new. We defer its details to Appendix A because it seems more instructive to present them independently. The other cases in (4.7)–(4.9) are all reduced to (4.6). For instance, assume that $\|SG_{\alpha,p}(A,B)\|=\|G_{\alpha,q}(A,B)\|$ for $\alpha,p,q$ satisfying one of the two conditions in (4.7). When $0<\alpha<1$ with $p,q>0$ arbitrary, for a $q^{\prime}\in(0,q)$ by Theorem 3.1(b) and Proposition 3.13(1) one has
$\|G_{\alpha,q}(A,B)\|\leq\|G_{\alpha,q^{\prime}}(A,B)\|\leq\|SG_{\alpha,p}(A,B)\|,$
implying that $\|G_{\alpha,q}(A,B)\|=\|G_{\alpha,q^{\prime}}(A,B)\|$ so that $AB=BA$ from the (4.6) case. When $1<\alpha\leq 2$ and $p/q<\max\bigl\{2,{\alpha\over\alpha-1}\bigr\}$, choose a $q^{\prime}\in(0,q)$ such that $\alpha,p,q^{\prime}$ satisfy the same condition. By Theorem 3.1(c) and Proposition 3.13(3) one has
$\|SG_{\alpha,p}(A,B)\|\leq\|G_{\alpha,q^{\prime}}(A,B)\|\leq\|G_{\alpha,q}(A,B)\|,$
implying that $\|G_{\alpha,q^{\prime}}(A,B)\|=\|G_{\alpha,q}(A,B)\|$ so that $AB=BA$. The proof is similar for (4.8) by using Proposition 3.14(2) and (3). For (4.9), (b) and (c) of Proposition 3.17 are used for $0<\alpha<1$ and $1<\alpha\leq 2$, respectively. ∎
The next proposition refines the log-majorizations given in (a), (b) and (c) of Theorem 3.1 into the local characterization in the sense that those log-majorizations are characterized for any fixed non-commuting $A,B>0$.
Proposition 4.2.
Let $p,q>0$ and $A,B>0$ be such that $AB\neq BA$.
  • (1)
    For any $\alpha\in(0,\infty)\setminus\{1\}$ we have
    $R_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)\iff\mathrm{Tr}\,R_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,q}(A,B)\iff p\leq q.$
    For $0<\alpha<1$ the above equivalences hold for $A,B\geq 0$ with $AB\neq BA$.
  • (2)
    For $0<\alpha<1$ we have
    $G_{\alpha,q}(A,B)\prec_{\log}G_{\alpha,p}(A,B)\iff\mathrm{Tr}\,G_{\alpha,q}(A,B)\leq\mathrm{Tr}\,G_{\alpha,p}(A,B)\iff p\leq q.$
  • (3)
    For $1<\alpha\leq 2$ we have
    $G_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,q}(A,B)\iff\mathrm{Tr}\,G_{\alpha,p}(A,B)\leq\mathrm{Tr}\,G_{\alpha,q}(A,B)\iff p\leq q.$
Proof.
(1) Let $A,B\geq 0$ when $0<\alpha<1$, while $A,B>0$ when $\alpha>1$. If $p\leq q$, then $R_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ by Theorem 3.1(a). Obviously, $R_{\alpha,p}(A,B)\prec_{\log}R_{\alpha,q}(A,B)$ implies $\mathrm{Tr}\,R_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,q}(A,B)$. Finally, assume that $\mathrm{Tr}\,R_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,q}(A,B)$. If $p>q$, then we must have $\mathrm{Tr}\,R_{\alpha,p}(A,B)=\mathrm{Tr}\,R_{\alpha,q}(A,B)$ by Theorem 3.1(a) again. This implies $AB=BA$ by case (4.1) of Theorem 4.1(1), contradicting the assumption $AB\neq BA$. Hence (1) follows.
(2) and (3) are proved similarly to (1) by use of Theorem 3.1(b), (c) and case (4.6) of Theorem 4.1(2) (i.e., Theorem A.1 in Appendix A). ∎
Remark 4.3.
It is not possible to have the local characterization similar to Proposition 4.2 for log-majorizaion between $G_{\alpha,p}$ and $R_{\alpha,q}$ for $\alpha>1$. In fact, log-majorization between $G_{\alpha,p}$ and $R_{\alpha,q}$ is indefinite when $\alpha>1$ and $\min\{\alpha/2,\alpha-1\}<p/q<\max\{\alpha/2,\alpha-1\}$; see Theorem 3.1(e) and (f).
5 Joint concavity/convexity
The aim of this section is to examine the joint concavity/convexity of the trace functions for quasi-geometric type matrix means of our concern.
5.1 Joint concavity/convexity and monotonicity
Before going to our main discussions of this section, we first review a known result on the joint concavity of trace functions for Kubo–Ando’s operator means and then a strong connection of the joint concavity/convexity and the monotonicity property of our target trace functions.
The next theorem was shown in [30, Theorem 3.2] in a more general form by a complex function method using Pick functions (called Epstein’s method).
Theorem 5.1.
Let $\sigma$ be a Kubo–Ando’s operator mean. If $0\leq p,q\leq 1$ and $0\leq s\leq 1/\max\{p,q\}$, then the function $(A,B)\mapsto\mathrm{Tr}(A^{p}\sigma B^{q})^{s}$ is jointly concave on $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ for any $n\geq 1$. In particular, if $0<p\leq 1$, then $(A,B)\mapsto\mathrm{Tr}(A^{p}\sigma B^{p})^{1/p}$ is jointly concave on $\mathbb{M}_{n}\times\mathbb{M}_{n}^{+}$.
In particular, $\mathrm{Tr}\,\mathcal{A}_{\alpha,p}$, $\mathrm{Tr}\,\mathcal{H}_{\alpha,p}$ and $\mathrm{Tr}\,G_{\alpha,p}$ are jointly concave for $0<\alpha\leq 1$ and $0<p\leq 1$. More specified results on the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{A}_{\alpha,p}$ and $\mathrm{Tr}\,\mathcal{H}_{\alpha,p}$ are known as follows (see [8, 15] and [30, p. 1583]):
Proposition 5.2.
Let $0<\alpha<1$, $p>0$ and $n\in\mathbb{N}$, $n\geq 2$ be arbitrary. Then the following hold:
  • (1)
    $(A,B)\mapsto\mathrm{Tr}\,\mathcal{A}_{\alpha,p}(A,B)$ is jointly concave on $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ if and only if $0<p\leq 1$.
  • (2)
    $(A,B)\mapsto\mathrm{Tr}\,\mathcal{A}_{\alpha,p}(A,B)$ is jointly convex on $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ if and only if $1\leq p\leq 2$.
  • (3)
    $(A,B)\mapsto\mathrm{Tr}\,\mathcal{H}_{\alpha,p}(A,B)$ is jointly concave on $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ if and only if $0<p\leq 1$.
  • (4)
    $(A,B)\mapsto\mathrm{Tr}\,\mathcal{H}_{\alpha,p}(A,B)$ is not jointly convex on $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ for any $p>0$.
Below we write
$(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}:=\{(A,B)\in\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}:s(A)\geq s(B)\}.$
Consider a function $Q:(A,B)\in\mathrm{dom}\,Q\mapsto Q(A,B)\in[0,+\infty)$ with the domain
$\mathrm{dom}\,Q=\bigsqcup_{n\in\mathbb{N}}(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})\quad\mbox{or}\quad\bigsqcup_{n\in\mathbb{N}}(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}.$ (5.1)
Note that $\mathrm{dom}_{n}Q:=\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ or $(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}$ is a convex cone, and if $(A_{1},B_{1})\in\mathrm{dom}_{n}Q$ and $(A_{2},B_{2})\in\mathrm{dom}_{m}Q$, then $(A_{1}\oplus A_{2},B_{1}\oplus B_{2})\in\mathrm{dom}_{n+m}Q$ and $(A_{1}\otimes A_{2},B_{1}\otimes B_{2})\in\mathrm{dom}_{nm}Q$. Consider the following properties of $Q$:
  • (a)
    Normalization identity: $Q(A,A)=\mathrm{Tr}\,A$ for all $A\geq 0$.
  • (b)
    Positive homogeneity: $Q(\lambda A,\lambda B)=\lambda Q(A,B)$ for all $(A,B)\in\mathrm{dom}\,Q$ and any $\lambda\geq 0$.
  • (c)
    Additivity under direct sums: $Q(A_{1}\oplus A_{2},B_{1}\oplus B_{2})=Q(A_{1},B_{1})+Q(A_{2},B_{2})$ for all $(A_{i},B_{i})\in\mathrm{dom}\,Q$, $i=1,2$.
  • (d)
    Multiplicativity under tensor products: $Q(A_{1}\otimes A_{2},B_{1}\otimes B_{2})=Q(A_{1},B_{1})Q(A_{2},B_{2})$ for all $(A_{i},B_{i})\in\mathrm{dom}\,Q$, $i=1,2$.
  • (e)
    Unitary invariance: $Q(UAU^{*},UBU^{*})=Q(A,B)$ for all $(A,B)\in\mathrm{dom}_{n}\,Q$ and all unitaries $U\in\mathbb{M}_{n}$, $n\in\mathbb{N}$.
A linear map $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$ is positive if $A\in\mathbb{M}_{n}^{+}$ implies $\Phi(A)\in\mathbb{M}_{m}^{+}$. It is said to be completely positive if for each $k\in\mathbb{N}$ the map $\mathrm{id}_{k}\otimes\Phi$ defined by $(\mathrm{id}_{k}\otimes\Phi)([A_{ij}]_{i,j=1}^{k})=[\Phi(A_{ij})]_{i,j=1}^{k}$ is positive. When $\Phi$ is completely positive and trace-preserving, i.e., $\mathrm{Tr}\,\Phi(A)=\mathrm{Tr}\,A$ for all $A\in\mathbb{M}_{n}$, $\Phi$ is called a CPTP map or a quantum channel. This is a very important notion in quantum information. Note that if $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$ is positive and $(A,B)\in(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}$, then $(\Phi(A),\Phi(B))\in(\mathbb{M}_{m}^{+}\times\mathbb{M}_{m}^{+})_{\geq}$.
The next theorem is rather well known to experts on quantum Rényi divergences, while we give a proof in Appendix B for the convenience of the reader.
Theorem 5.3.
Let $Q:\mathrm{dom}\,Q\to[0,+\infty)$ be as given above with the domain in (5.1). Consider the following conditions:
  • (i)
    $Q(\Phi(A),\Phi(B))\geq Q(A,B)$ for all $(A,B)\in\mathrm{dom}_{n}\,Q$ and all CPTP maps $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$, $n,m\in\mathbb{N}$.
  • (i )
    $Q(\Phi(A),\Phi(B))\leq Q(A,B)$ for all $(A,B)\in\mathrm{dom}_{n}\,Q$ and all CPTP maps $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$, $n,m\in\mathbb{N}$.
  • (ii)
    $Q$ is jointly concave on $\mathrm{dom}_{n}\,Q$ for all $n\in\mathbb{N}$.
  • (ii )
    $Q$ is jointly convex on $\mathrm{dom}_{n}\,Q$ for all $n\in\mathbb{N}$.
Then the following hold:
  • (1)
[…]
  • (2)
[…]
  • (3)
[…]
5.2 Quantum Rényi type divergences
In this subsection we briefly review quantum $\alpha$-Rényi type divergences from quantum information and present certain necessary conditions for the joint concavity/convexity of our quasi-geometric type matrix means (see Theorem 5.8 below).
Definition 5.4.
Let $\alpha\in(0,\infty)\setminus\{1\}$. The classical $\alpha$-Rényi divergence $D_{\alpha}^{\mathrm{cl}}(b\|a)$ is defined for non-negative vectors $a,b\in[0,\infty)^{n}$ with $b\neq 0$ by
$D_{\alpha}^{\mathrm{cl}}(b\|a):=\begin{cases}{1\over\alpha-1}\log{\sum_{i=1}^{n}b_{i}^{\alpha}a_{i}^{1-\alpha}\over\sum_{i=1}^{n}b_{i}}&\text{if $0<\alpha<1$ or $\mathrm{supp}\,a\supset\mathrm{supp}\,b$},\\ +\infty&\text{if $\alpha>1$ and $\mathrm{supp}\,a\not\supset\mathrm{supp}\,b$},\end{cases}$
where $\mathrm{supp}\,a:=\{i:a_{i}>0\}$ and $0^{1-\alpha}:=0$ for $\alpha>1$. We say that a function
$D_{\alpha}^{\mathrm{q}}:(A,B)\in\bigsqcup_{n\in\mathbb{N}}(\mathbb{M}_{n}^{+}\times(\mathbb{M}_{n}^{+}\setminus\{0\}))\mapsto D_{\alpha}^{\mathrm{q}}(B\|A)\in(-\infty,+\infty]$
is a quantum $\alpha$-Rényi type divergence if it is invariant under isometries, i.e.,
$D_{\alpha}^{\mathrm{q}}(VBV^{*}\|VAV^{*})=D_{\alpha}^{\mathrm{q}}(B\|A),\qquad A,B\in\mathbb{M}_{n}^{+},\ B\neq 0,\ n\in\mathbb{N}$
for any isometry $V:\mathbb{C}^{n}\to\mathbb{C}^{m}$ and satisfies
$D_{\alpha}^{\mathrm{q}}(\mathrm{diag}(b)\|\mathrm{diag}(a))=D_{\alpha}^{\mathrm{cl}}(b\|a),\qquad a,b\in[0,\infty)^{n},\ b\neq 0,n\in\mathbb{N},$
where $\mathrm{diag}(a)$ is the diagonal matrix whose diagonals are entries of $a$.
When $A,B\in\mathbb{M}_{n}^{+}$ are commuting, one can write $A=\sum_{i=1}^{n}a_{i}|e_{i}\rangle\langle e_{i}|$ and $B=\sum_{i=1}^{n}b_{i}|e_{i}\rangle\langle e_{i}|$ for some orthonormal basis $(e_{i})_{i=1}^{n}$ of $\mathbb{C}^{n}$ and by unitary invariance one has
$D_{\alpha}^{q}(B\|A)=D_{\alpha}^{q}(\mathrm{diag}(b)\|\mathrm{diag}(a))=D_{\alpha}^{\mathrm{cl}}(b\|a).$
In view of this, we may write $D_{\alpha}^{q}(B\|A)=D_{\alpha}^{\mathrm{cl}}(B\|A)$ when $AB=BA$. There are the notions of the minimal and the maximal quantum $\alpha$-Rényi divergences. The maximal one is Matsumoto’s maximal $\alpha$-Rényi divergence [44] defined by
$D_{\alpha}^{\max}(B\|A):=\inf\{D_{\alpha}^{\mathrm{cl}}(b\|a):\Gamma(a)=A,\ \Gamma(b)=B\},$ (5.2)
where the infimum is taken over triplets $(\Gamma,a,b)$ (called reverse tests) consisting of $a,b\in[0,\infty)^{m}$ and a (completely) positive trace-preserving map $\Gamma:\mathbb{C}^{m}\to\mathbb{M}_{n}$ satisfying $\Gamma(a)=A$, $\Gamma(b)=B$ for some $m\in\mathbb{N}$. The minimal one is the so-called measured $\alpha$-Rényi divergence defined for $A,B\in\mathbb{M}_{n}^{+}$ by
$D_{\alpha}^{\mathrm{meas}}(B\|A):=\sup\bigl\{D_{\alpha}^{\mathrm{cl}}\bigl((\mathrm{Tr}\,M_{i}B)_{i=1}^{k}\big\|(\mathrm{Tr}\,M_{i}A)_{i=1}^{k}\bigr):\mbox{$(M_{i})_{i=1}^{k}$ is a POVM on $\mathbb{C}^{n}$, $k\in\mathbb{N}$}\bigr\},$ (5.3)
where a POVM on $\mathbb{C}^{n}$ is a family $(M_{i})_{i=1}^{k}\subset\mathbb{M}_{n}^{+}$ with $\sum_{i=1}^{k}=I_{n}$ (i.e., a $k$-outcome positive operator-valued measure). Furthermore, the regularized measured $\alpha$-Rényi divergence is defined by
$\overline{D}_{\alpha}^{\mathrm{meas}}(B\|A):=\sup_{m\geq 1}{1\over m}\,D_{\alpha}^{\mathrm{meas}}(B^{\otimes m}\|A^{\otimes m})=\lim_{m\to\infty}{1\over m}\,D_{\alpha}^{\mathrm{meas}}(B^{\otimes m}\|A^{\otimes m}),$ (5.4)
where the last equality holds since $m\in\mathbb{N}\mapsto D_{\alpha}^{\mathrm{meas}}(B^{\otimes m}\|A^{\otimes m})$ is a superadditive sequence.
Now assume that a quantum $\alpha$-Rényi type divergence $D_{\alpha}^{\mathrm{q}}$ satisfies the monotonicity under CPTP maps, i.e.,
$D_{\alpha}^{\mathrm{q}}(\Phi(B)\|\Phi(A))\leq D_{\alpha}^{\mathrm{q}}(B\|A),\qquad A,B\geq 0$
for any CPTP map $\Phi$. Then from definitions (5.2) and (5.3) it follows that
$D_{\alpha}^{\mathrm{meas}}(B\|A)\leq D_{\alpha}^{\mathrm{q}}(B\|A)\leq D_{\alpha}^{\max}(B\|A),\qquad A,B\geq 0.$ (5.5)
Indeed, for any POVM $(M_{i})_{i=1}^{k}$ on $\mathbb{C}^{n}$, since $\Phi(A):=\mathrm{diag}((\mathrm{Tr}\,M_{i}A)_{i=1}^{k})$ for $A\in\mathbb{M}_{n}$ is a CPTP map, we have
$D_{\alpha}^{\mathrm{cl}}\bigl((\mathrm{Tr}\,M_{i}B)_{i=1}^{k}\big\|(\mathrm{Tr}\,M_{i}A)_{i=1}^{k}\bigr)=D_{\alpha}^{\mathrm{q}}\bigl(\mathrm{diag}((\mathrm{Tr}\,M_{i}B)_{i=1}^{k})\big\|\mathrm{diag}((\mathrm{Tr}\,M_{i}A)_{i=1}^{k})\bigr)\leq D_{\alpha}^{\mathrm{q}}(B\|A),$
so that the first inequality in (5.5) follows from (5.3). On the other hand, for any reverse test $(\Gamma,a,b)$ with $\Gamma:\mathbb{C}^{m}\to\mathbb{M}_{n}$ and $a,b\in[0,\infty)^{m}$, define $\mathcal{E}:\mathbb{M}_{m}\to\mathbb{C}^{m}$ by $A=[a_{ij}]_{i,j=1}^{m}\mapsto(a_{11},\dots,a_{mm})$ and $\Phi:=\Gamma\circ\mathcal{E}:\mathbb{M}_{m}\to\mathbb{M}_{n}$; then $\Phi$ is a CPTP map and $\Phi(\mathrm{diag}(a))=\Gamma(a)$, $a\in\mathbb{C}^{m}$. Since
$D_{\alpha}^{\mathrm{q}}(B\|A)=D_{\alpha}^{\mathrm{q}}(\Gamma(b)\|\Gamma(a))=D_{\alpha}^{\mathrm{q}}(\Phi(\mathrm{diag}(b))\|\Phi(\mathrm{diag}(a))\bigr)\leq D_{\alpha}^{\mathrm{q}}(\mathrm{diag}(b)\|\mathrm{diag}(a))=D_{\alpha}^{\mathrm{cl}}(b\|a),$
we have the second inequality in (5.5). If, in addition, $D_{\alpha}^{\mathrm{q}}$ is additive under tensor products, then by definition (5.4) as well as (5.5) we furthermore have
$\overline{D}_{\alpha}^{\mathrm{meas}}(B\|A)\leq D_{\alpha}^{\mathrm{q}}(B\|A),\qquad A,B\geq 0.$ (5.6)
The so-called $\alpha$-$z$-Rényi divergence $D_{\alpha,z}$ [5] for $\alpha\in(0,\infty)\setminus\{1\}$ and $z>0$ is typical among quantum $\alpha$-Rényi type divergences. Their most important special cases are the Petz type $\alpha$-Rényi divergence $D_{\alpha}=D_{\alpha,1}$ [50] for $z=1$ and the sandwiched $\alpha$-Rényi divergence $\widetilde{D}_{\alpha}=D_{\alpha,\alpha}$ [49, 52] for $z=\alpha$. See, e.g., [45] for further details on quantum $\alpha$-Rényi divergences.
One can easily check the next lemma, whose proofs may be omitted.
Lemma 5.5.
Let $\mathcal{M}_{\alpha,p}$ be any of $R_{\alpha,p}$, $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$ with any $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$. Then $Q:=\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ satisfies all properties of (a)–(e) above.
For a quasi matrix mean $\mathcal{M}_{\alpha,p}$ as in Lemma 5.5 it is meaningful to associate the quantum divergence of Rényi type with two parameters $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$ as follows: for $A,B\in\mathbb{M}_{n}^{+}$ with $B\neq 0$,
$D^{\mathcal{M}_{\alpha,p}}(B\|A):=\begin{cases}{1\over\alpha-1}\log{\mathrm{Tr}\,\mathcal{M}_{\alpha,p}(A,B)\over\mathrm{Tr}\,B}&\text{if $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$},\\ +\infty&\text{otherwise},\end{cases}$ (5.7)
where $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$ has been fixed in Section 2 (see the paragraph after (2.8)). Here, note that the orders of the variables $A,B$ in $\mathcal{M}_{\alpha,p}(A,B)$ and $D^{\mathcal{M}_{\alpha,p}}(B\|A)$ are opposite; we follow the convention of Kubo–Ando’s operator means for the former and the usual convention of quantum divergences for the latter. The next lemma is also easy to verify.
Lemma 5.6.
Let $p>0$. Either when $\mathcal{M}\in\{R,G,LG\}$ and $\alpha\in(0,\infty)\setminus\{1\}$, or when $\mathcal{M}\in\{SG,\widetilde{S}G\}$ and $\alpha>1$, the function $D^{\mathcal{M}_{\alpha,p}}$ defined in (5.7) is a quantum $\alpha$-Rényi type divergence in the sense of Definition 5.4, and is additive under tensor products.
When $\mathcal{M}\in\{SG,\widetilde{S}G\}$ and $0<\alpha<1$, $D^{\mathcal{M}_{\alpha,p}}$ is invariant under isometries and satisfies $D^{\mathcal{M}_{\alpha,p}}(\mathrm{diag}(b)\|\mathrm{diag}(a))=D_{\alpha}^{\mathrm{cl}}(b\|a)$ for all $a,b\in[0,\infty)$, $b\neq 0$, with $\mathrm{supp}\,a\supset\mathrm{supp}\,b$, and it is additive under tensor product.
We note that $D^{R_{\alpha,1/z}}$ becomes the $\alpha$-$z$-Rényi divergence $D_{\alpha,z}$; in particular, $D^{R_{\alpha,1}}$ is the Petz type $D_{\alpha}$ and $D^{R_{\alpha,1/\alpha}}$ is the sandwiched $\widetilde{D}_{\alpha}$. Also, $D^{G_{\alpha,1}}$ is the maximal $f$-divergence [35] associated to the function $f(x):=x^{\alpha}$ ($x>0$). Here, in the next theorem we recall a few results related to the minimal and the maximal bounds in (5.5) and (5.6), as we will use those below.
Theorem 5.7.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $A,B\geq 0$, $B\neq 0$.
  • (1)
    ([44]) When $0<\alpha\leq 2$, $D_{\alpha}^{\max}(B\|A)=D^{G_{\alpha,1}}(B\|A)$.
  • (2)
    ([45, Remark II.6]) When $\alpha>2$, $D_{\alpha}^{\max}(B\|A)\leq D^{G_{\alpha,1}}(B\|A)$ and $D_{\alpha}^{\max}\neq D^{G_{\alpha,1}}$.
  • (3)
    ([46, Theorem 3.7], [27]) When $1/2\leq\alpha<\infty$,
    $\overline{D}_{\alpha}^{\mathrm{meas}}(B\|A)=\widetilde{D}_{\alpha}(B\|A)=\lim_{m\to\infty}{1\over m}D_{\alpha}^{\mathrm{cl}}(\mathcal{E}_{A^{\otimes m}}(B^{\otimes m})\|A^{\otimes m}),$
    where $\mathcal{E}_{A^{\otimes m}}$ is the pinching with respect to $A^{\otimes m}$.
  • (4)
    ([9, Theorem 7]) When $0<\alpha<1/2$, $\widetilde{D}_{\alpha}(B\|A)\leq D_{\alpha}^{\mathrm{meas}}(B\|A)\leq\overline{D}_{\alpha}^{\mathrm{meas}}(B\|A)$, and $\widetilde{D}_{\alpha}(B\|A)<D_{\alpha}^{\mathrm{meas}}(B\|A)$ if $AB\neq BA$.
For any $\mathcal{M}_{\alpha,p}$ in Lemma 5.5, since $\mathcal{M}_{\alpha,p}(a,b)=a^{1-\alpha}b^{\alpha}$ in $a,b\in(0,\infty)$ is not convex when $0<\alpha<1$ and not concave when $\alpha>1$, we may consider only the joint concavity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ on $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ for $0<\alpha<1$, and the joint convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ on ($\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}$ when $\alpha>1$. Below we will simply write $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ is jointly concave/convex” if it is so on its domain $\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+}$ or $(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}$ for any $n\in\mathbb{N}$, equivalently, if it is so on $\mathbb{M}_{n}^{++}\times\mathbb{M}_{n}^{++}$ for any $n\in\mathbb{N}$ (thanks to Proposition 2.1(2)).
The next theorem will repeatedly be used in the rest of this section.
Theorem 5.8.
Let $\mathcal{M}_{\alpha,p}$ be as in Lemma 5.5.
  • (i)
    If $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ is jointly concave, then $0<\alpha<1$ and
    $\mathrm{Tr}\,G_{\alpha,1}(A,B)\leq\mathrm{Tr}\,\mathcal{M}_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$
    for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$.
  • (ii)
    If $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ is jointly convex, then $\alpha>1$ and
    $\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)\leq\mathrm{Tr}\,\mathcal{M}_{\alpha,p}(A,B)\leq\mathrm{Tr}\,G_{\alpha,1}(A,B)$
    for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$
Proof.
Assume that $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ is jointly concave as in (i), or jointly convex as in (ii). From the remark mentioned just above the theorem, the parameter $\alpha$ is restricted to $0<\alpha<1$ in (i) and to $\alpha>1$ in (ii). Unless $\mathcal{M}\in\{SG,\widetilde{S}G\}$ and $0<\alpha<1$, it follows from Lemmas 5.5, 5.6 and Theorem 5.3 that $D^{\mathcal{M}_{\alpha,p}}$ is a quantum $\alpha$-Rényi type divergence satisfying the monotonicity under CPTP maps and it is moreover additive under tensor products. Therefore, we can apply (5.5) and (5.6) to $D^{\mathcal{M}_{\alpha,p}}$ to obtain
$\overline{D}_{\alpha}^{\mathrm{meas}}(B\|A)\leq D^{\mathcal{M}_{\alpha,p}}(B\|A)\leq D_{\alpha}^{\max}(B\|A)$ (5.8)
for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$, $B\neq 0$. When $\mathcal{M}\in\{SG,\widetilde{S}G\}$ and $0<\alpha<1$, even though $D^{\mathcal{M}_{\alpha,p}}$ is not precisely a quantum $\alpha$-Rényi type divergence in Definition 5.4, we can still have (5.8) for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. In fact, Theorem 5.3 implies that $D^{\mathcal{M}_{\alpha,p}}$ is monotone under CPTP maps with restriction to $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. Hence by Lemma 5.6 the first inequality in (5.8) holds on $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$. Also, note (see [44] and the proof of [35, Proposition 4.1]) that for any $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$ (hence $s(A)\geq s(B)$) the infimum in (5.2) is attained by a reverse test (called the “minimal reverse test”) $(\Gamma,a,b)$ with $\mathrm{supp}\,a\supset\mathrm{supp}\,b$. Hence the second inequality in (5.8) holds on $\mathrm{dom}\,\mathcal{M}_{\alpha,p}$.
From (5.8) and (1)–(4) of Theorem 5.7 altogether we have
$D^{R_{\alpha,1/\alpha}}(B\|A)=\widetilde{D}_{\alpha}(B\|A)\leq D^{\mathcal{M}_{\alpha,p}}(B\|A)\leq D^{G_{\alpha,1}}(B\|A)$
for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$, $B\neq 0$. In view of (5.7) this shows the inequalities in (i) ($0<\alpha<1$) and in (ii) ($\alpha>1$) for all $(A,B)\in\mathrm{dom}\,\mathcal{M}_{\alpha,p}$ except for the case $B=0$. But note that $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}(A,0)=0$ for all $A\geq 0$ and all $\mathcal{M}_{\alpha,p}$’s. ∎
5.3 Joint concavity/convexity of trace functions
In this subsection we examine the joint concavity/convexity of $\mathrm{Tr}\,R_{\alpha,p}$, $\mathrm{Tr}\,LE_{\alpha}$, $\mathrm{Tr}\,G_{\alpha,p}$, $\mathrm{Tr}\,SG_{\alpha,p}$ and $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ with the domains described in Section 2.
The question of the joint concavity/convexity of $\mathrm{Tr}\,R_{\alpha,p}$ was an intriguing issue because it is equivalent to the monotonicity under CPTP maps of $D_{\alpha,z}$ (see Theorem 5.3). It was finally settled by Zhang [53] as follows:
Theorem 5.9 ([53]).
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$.
  • (1)
    $\mathrm{Tr}\,R_{\alpha,p}$ is jointly concave if and only if $0<\alpha<1$ and $1/p\geq\max\{\alpha,1-\alpha\}$.
  • (2)
    $\mathrm{Tr}\,R_{\alpha,p}$ is jointly convex if and only if $\alpha>1$ and $\max\{\alpha/2,\alpha-1\}\leq 1/p\leq\alpha$.
It is worthwhile to compare condition of Theorem 5.9(2) with the $p=1$ case of (iii) of Theorem 3.2(2). Notice that the former condition is strictly stronger than the latter though they are quite similar.
Remark 5.10.
Note that the ‘only if’ parts of (1) and (2) of Theorem 5.9 are verified by Theorem 5.8. Indeed, if $\mathrm{Tr}\,R_{\alpha,p}$ is jointly concave, then we have $0<\alpha<1$ and $1/p\geq\alpha$ by Theorem 5.8 and Proposition 4.2(1), so we have $1/p\geq 1-\alpha$ too since $R_{\alpha,p}(A,B)=R_{1-\alpha,p}(B,A)$. If $\mathrm{Tr}\,R_{\alpha,p}$ is jointly convex, then we have $\alpha>1$ and $\max\{\alpha/2,\alpha-1\}\leq 1/p\leq\alpha$.
Remark 5.11.
Although Theorem 5.9 implies the monotonicity of $D_{\alpha,z}$ under CPTP maps for $z=1/p$ satisfying the conditions in (1) and (2), it is in fact known [38, 34] (even in the von Neumann algebra setting) that $D_{\alpha,z}$ satisfies the monotonicity under general positive (not necessarily completely positive) trace-preserving maps for the same $z$.
The next theorem is concerned with the joint concavity/convexity of $\mathrm{Tr}\,LE_{\alpha}$.
Theorem 5.12.
Let $\alpha$ be as above.
  • (1)
    For any $\alpha\in(0,1)$, $\mathrm{Tr}\,LE_{\alpha}$ is jointly concave.
  • (2)
    For any $\alpha>1$, $\mathrm{Tr}\,LE_{\alpha}$ is not jointly convex on $\mathbb{M}_{2}^{++}\times\mathbb{M}_{2}^{++}$.
Proof.
(1) It follows from Theorem 5.9(1) that $\mathrm{Tr}\,R_{\alpha,p}$ is jointly concave for all $p>0$ with $1/p\geq\max\{\alpha,1-\alpha\}$. Letting $p\searrow 0$ shows the assertion due to Theorem 2.2.
(2) Let $\alpha>1$. Let $A:=\begin{bmatrix}a&0\\ 0&b\end{bmatrix}$ with $a,b>0$, $a\neq b$, and $B:={1\over 2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$, and let $\mathcal{E}_{A}$ be the pinching with respect to $A$. It was shown in [47, Lemma 3.17] that $\mathrm{Tr}\,LE_{\alpha}(A,\mathcal{E}_{A}(B))>\mathrm{Tr}\,LE_{\alpha}(A,B)$ for some choice of $a,b$. (Note that $\mathrm{Tr}\,LE_{\alpha}(A,B)$ is denoted as $Q_{\alpha}^{\flat}(B\|A)$ in [47].) Let $U:=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$. Since $A={1\over 2}(A+UAU^{*})$, $\mathcal{E}_{A}(B)={1\over 2}(B+UBU^{*})$ and $\mathrm{Tr}\,LE_{\alpha}(A,B)=\mathrm{Tr}\,LE_{\alpha}(UAU^{*},UBU^{*})$, we have
$\mathrm{Tr}\,LE_{\alpha}\Bigl({A+UAU^{*}\over 2},{B+UBU^{*}\over 2}\Bigr)>{\mathrm{Tr}\,LE_{\alpha}(A,B)+\mathrm{Tr}\,LE_{\alpha}(UAU^{*},UBU^{*})\over 2},$
showing the result. ∎
Remark 5.13.
In [47, Theorem 3.6] Mosonyi and Ogawa showed the variational expression
$\mathrm{Tr}\,LE_{\alpha}(A,B)=\max_{X\in\mathbb{M}_{n}^{+},s(X)\leq s(B)}\{\mathrm{Tr}\,X-(1-\alpha)D(X\|A)-\alpha D(X\|B)\}$ (5.9)
for any $\alpha\in(0,\infty)\setminus\{1\}$ and for every $A,B\in\mathbb{M}_{n}^{+}$, where $D$ is the Umegaki relative entropy. Thanks to the well known joint convexity of $D$, expression (5.9) implies the joint concavity in Theorem 5.12(1). Moreover, since $D$ satisfies the monotonicity under general positive trace-preserving maps (see [7, 48]), it easily follows from (5.9) that for every $\alpha\in(0,1)$ and for any positive trace-preserving map $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$,
$\mathrm{Tr}\,LE_{\alpha}(\Phi(A),\Phi(B))\geq\mathrm{Tr}\,LE_{\alpha}(A,B),\qquad A,B\in\mathbb{M}_{n}^{+}.$ (5.10)
From Remark 5.11 the same inequality holds for $\mathrm{Tr}\,R_{\alpha,p}$ if $1/p\geq\max\{\alpha,1-\alpha\}$. Letting $p\searrow 0$ as in the proof of Theorem 5.12(1) gives (5.10) as well.
Remark 5.14.
For $\alpha>1$, since $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,LE_{\alpha}$ fails to hold (see Proposition 3.17(a) and Remark 3.16(2)), we also confirm by Theorem 5.8(ii) that $\mathrm{Tr}\,LE_{\alpha}$ is not jointly convex.
The next theorem determines when $\mathrm{Tr}\,G_{\alpha,p}$ is jointly concave.
Theorem 5.15.
Let $\alpha,p$ be as in Theorem 5.9. Then $\mathrm{Tr}\,G_{\alpha,p}$ is jointly concave if and only if $0<\alpha<1$ and $p\leq 1$.
Proof.
If $0<\alpha<1$ and $0<p\leq 1$, then Theorem 5.1 implies that $\mathrm{Tr}\,G_{\alpha,p}$ is jointly concave. Conversely, assume the joint concavity of $\mathrm{Tr}\,G_{\alpha,p}$. Then by Theorem 5.8(i) we have $0<\alpha<1$ and $\mathrm{Tr}\,G_{\alpha,1}\leq\mathrm{Tr}\,G_{\alpha,p}$, which yields $p\leq 1$ by Proposition 4.2(2). ∎
For $\alpha>1$ the complete characterization of the joint convexity of $\mathrm{Tr}\,G_{\alpha,p}$ is not known, while we give some partial results in the following:
Proposition 5.16.
Let $\alpha>1$ and $p>0$.
  • (1)
    Assume that $\mathrm{Tr}\,G_{\alpha,p}$ is jointly convex. Then we have $\max\bigl\{1/2,{\alpha-1\over\alpha}\bigr\}\leq p\leq 1$.
  • (2)
    For $\alpha=2$, $\mathrm{Tr}\,G_{2,p}$ is jointly convex if and only if $1/2\leq p\leq 1$.
  • (3)
    For $p=1$, $\mathrm{Tr}\,G_{\alpha,1}$ is jointly convex if and only if $1<\alpha\leq 2$.
  • (4)
    If $1<\alpha\leq 2$ and $0<p\leq 1$, then for any fixed $B\in\mathbb{M}_{n}^{+}$, $A\mapsto\mathrm{Tr}\,G_{\alpha,p}(A,B)$ is convex on $\{A\in\mathbb{M}_{n}^{+}:s(A)\geq s(B)\}$, $n\in\mathbb{N}$.
Proof.
(1) By Theorem 5.8(ii) the assumption implies that
$\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)\leq\mathrm{Tr}\,G_{\alpha,p}(A,B)\leq\mathrm{Tr}\,G_{\alpha,1}(A,B)$
for all $A,B\in\mathbb{M}_{n}^{++}$, $n\in\mathbb{N}$. This in particular yields that
$R_{\alpha,1/\alpha}(A,B)\prec_{\log}G_{\alpha,p}(A,B)\prec_{\log}G_{\alpha,1}(A,B),\qquad A,B\in\mathbb{M}_{2}^{++}.$
From the above two log-majorizations we must have $\alpha p\geq\max\{\alpha/2,\alpha-1\}$ by Theorem 3.2(2) and $p\leq 1$ by Proposition 3.11(1). Hence the result follows.
(2) Since $\mathrm{Tr}\,G_{2,p}=\mathrm{Tr}\,R_{2,p}$ (Remark 2.4(3)), we see by Theorem 5.9(2) that $\mathrm{Tr}\,G_{2,p}$ is jointly convex if and only if $1/2\leq p\leq 1$.
(3) The ‘if’ part is a general result for the operator perspective for operator convex functions on $(0,\infty)$ due to [18, 19]. For the ‘only if’ part, a more general and stronger result was shown in [35, Proposition A.1]; see Remark 5.17 below.
(4) In view of Proposition 2.1(2) it suffices to show that for any fixed $B\in\mathbb{M}_{n}^{++}$, $A\mapsto\mathrm{Tr}\,G_{\alpha,p}(A,B)$ is convex on $\mathbb{M}_{n}^{++}$. For any $A,B\in\mathbb{M}_{n}^{++}$ we can write
$\mathrm{Tr}\,G_{\alpha,p}(A,B)=\big\|B^{p/2}(B^{-p/2}A^{p}B^{-p/2})^{1-\alpha}B^{p/2}\big\|_{1/p}^{1/p}.$
Now let $A_{1},A_{2},B\in\mathbb{M}_{n}^{++}$ and $0<\lambda<1$. Since $(\lambda A_{1}+(1-\lambda)A_{2})^{p}\geq\lambda A_{1}^{p}+(1-\lambda)A_{2}^{p}$ thanks to $0<p\leq 1$, one has
$B^{-p/2}(\lambda A_{1}+(1-\lambda)A_{2})^{p}B^{-p/2}\geq\lambda B^{-p/2}A_{1}^{p}B^{-p/2}+(1-\lambda)B^{-p/2}A_{2}^{p}B^{-p/2}.$
Since $1<1-\alpha<0$, note that $x^{1-\alpha}$ ($x>0$) is operator monotone decreasing and operator convex. Hence one has
$B^{p/2}\bigl(B^{-p/2}(\lambda A_{1}+(1-\lambda)A_{2})^{p}B^{-p/2}\bigr)^{1-\alpha}B^{p/2}$
$\quad\leq B^{p/2}\bigl(\lambda B^{-p/2}A_{1}^{p}B^{-p/2}+(1-\lambda)B^{-p/2}A_{2}^{p}B^{-p/2}\bigr)^{1-\alpha}B^{p/2}$
$\quad\leq\lambda B^{p/2}\bigl(B^{-p/2}A_{1}^{p}B^{-p/2}\bigr)^{1-\alpha}B^{p/2}+(1-\lambda)B^{p/2}\bigl(B^{-p/2}A_{2}^{p}B^{-p/2}\bigr)^{1-\alpha}B^{p/2}.$
Since $\|\cdot\|_{1/p}^{1/p}$ is convex on $\mathbb{M}_{n}^{+}$, it follows that
$\mathrm{Tr}\,G_{\alpha,p}(\lambda A_{1}+(1-\lambda)A_{2},B)\leq\lambda\mathrm{Tr}\,G_{\alpha,p}(A_{1},B)+(1-\lambda)\mathrm{Tr}\,G_{\alpha,p}(A_{2},B),$
as desired. ∎
Remark 5.17.
In view of the proof of [35, Proposition A.1] together with [27, Proposition 3.1], Proposition 5.16(3) is in fact improved in such a way that either if $A\mapsto\mathrm{Tr}\,G_{\alpha,1}(A,B)$ on $\mathbb{M}_{2}^{++}$ for any fixed $B\in\mathbb{M}_{2}^{++}$ or if $B\mapsto\mathrm{Tr}\,G_{\alpha,1}(A,B)$ on $\mathbb{M}_{2}^{++}$ for any fixed $A\in\mathbb{M}_{2}^{++}$ is convex, then $1<\alpha\leq 2$.
Problem 5.18.
No sufficient condition for the joint convexity of $\mathrm{Tr}\,G_{\alpha,p}$ is known except the particular cases in (2) and (3) of Proposition 5.16. It is of our special concern whether or not $\mathrm{Tr}\,G_{\alpha,p}$ is jointly convex for $1<\alpha<2$ and $1/2\leq p<1$ (see remark (5) in Section 6).
As for $\mathrm{Tr}\,SG_{\alpha,p}$ and $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ we have the following:
Proposition 5.19.
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$.
  • (1)
    Let $0<\alpha<1$. If $\mathrm{Tr}\,SG_{\alpha,p}$ is jointly concave, then $p\leq\min\bigl\{{1-\alpha\over\alpha},{\alpha\over 1-\alpha}\bigr\}$. If $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ is jointly concave, then $p\leq 1$.
  • (2)
    For any $\alpha>1$ and any $p>0$, neither $\mathrm{Tr}\,SG_{\alpha,p}$ nor $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ is jointly convex.
Proof.
(1) For $0<\alpha<1$, if $\mathrm{Tr}\,SG_{\alpha,p}$ is jointly concave, then by Theorem 5.8(i),
$\mathrm{Tr}\,G_{\alpha,1}(A,B)\leq\mathrm{Tr}\,SG_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$
for all $A,B\geq 0$ with $s(A)\geq s(B)$. The first inequality gives no restriction (see Proposition 3.13(1)) but the second does the restriction $\alpha p\leq\min\{\alpha,1-\alpha\}$ so that $p\leq\min\bigl\{1,{1-\alpha\over\alpha}\bigr\}$. One can replace $\alpha$ with $1-\alpha$ to have $p\leq\min\bigl\{1,{\alpha\over 1-\alpha}\}$. Therefore, $p\leq\min\bigl\{{1-\alpha\over\alpha},{\alpha\over 1-\alpha}\bigr\}$ must hold. Next, if $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ is jointly concave, then $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}(A,B)\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$, which gives $p\leq 1$ by Theorem 3.6(1).
(2) For $\alpha>1$, assume that $\mathrm{Tr}\,SG_{\alpha,p}$ is jointly convex; then by Theorem 5.8(ii), $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,SG_{\alpha,p}$ on $\mathbb{M}_{2}^{++}\times\mathbb{M}_{2}^{++}$ must hold. But this contradicts Theorem 3.4(2). Assume that $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ is jointly convex. Then we must have $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ on $\mathbb{M}_{2}^{++}\times\mathbb{M}_{2}^{++}$, which contradicts Theorem 3.7(1). ∎
Problem 5.20.
We have no sufficient condition for the joint concavity of $\mathrm{Tr}\,SG_{\alpha,p}$ or $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$, except for the particular case $\alpha=1/2$, the case that $\mathrm{Tr}\,SG_{\alpha,p}=\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}=\mathrm{Tr}\,R_{1/2,2p}$.
6 Concluding remarks
(1) In the present paper we have followed the previous paper [33] to discuss the $\alpha$-weighted quasi-geometric type matrix means $R_{\alpha,p}$, $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$ not only for $0<\alpha<1$ but also $\alpha>1$ as defined in Section 2. We are concerned with log-majorizations $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for pairs $(\mathcal{M},\mathcal{N})$ from $R,G,SG,\widetilde{S}G,LE$ and for $\alpha\in(0,\infty)\setminus\{1\}$, $p>0$. Already known cases are summarized in Theorem 3.1. In Sections 3.13.4 we have examined unknown cases with the aim of finding the necessary and sufficient condition on $p,q,\alpha$ under which $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ holds for all $A,B>0$. But in many cases we have only a certain necessary condition and/or a certain sufficient condition for that, and the problem is still left open as explained in Problems 3.5, 3.8, 3.12 and 3.20. Especially interesting for us is to characterize $SG_{\alpha,p}\prec_{\log}SG_{\alpha,q}$ and $\widetilde{S}G_{\alpha,p}\prec_{\log}\widetilde{S}G_{\alpha,q}$ (see Problem 3.12).
(2) Another topic we have addressed is the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for each $\mathcal{M}$ from $R,G,SG,\widetilde{S}G,LE$. The cases of $\mathrm{Tr}\,R_{\alpha,p}$ ($\alpha>0$) and $\mathrm{Tr}\,G_{\alpha,p}$ ($0<\alpha<1$) are already known [53, 30]. In Section 5 we have aimed at characterizing the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for unknown cases. Our method is based on the close connection between the joint concavity/convexity of geometric type trace functions and the monotonicity under CPTP maps of the generated Rényi type divergences described in Theorem 5.3. Appealing to theory of quantum $\alpha$-Rényi divergences we have obtained a necessary condition for $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ to be jointly concave/convex. The condition is the sandwiched inequalities of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ with $\mathrm{Tr}\,R_{\alpha,1/\alpha}$ and $\mathrm{Tr}\,G_{\alpha,1}$ (see Theorem 5.8), which are related to the log-majorizations discussed in Section 3. However, we cannot obtain a sufficient condition for the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ in this way, while the necessary condition obtained can be best possible, for example, as exemplified in Remark 5.10. Thus the problem is still left open for $\mathrm{Tr}\,G_{\alpha,p}$ ($\alpha>1$), $\mathrm{Tr}\,SG_{\alpha,p}$ and $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ as mentioned in Problems 5.18 and 5.20.
(3) We have used the Lie–Trotter–Kato product formula (Theorem 2.2 and [33, Theorem 2.3]) in the proofs of (1), (4) and (5) of Corollary 3.18 and Theorem 5.12(1) for example. In fact, in view of Proposition 2.1(2) we may assume that $A,B>0$ in the proofs of those, in which case the proof of Theorem 2.2 becomes much simpler. Note however that in the proof of Proposition 2.1(2) we have used [33, Lemma A.3] that is essential in the proof of [33, Theorem 2.3].
(4) Appendix C below is a supplement to Theorems 5.3 and 5.8. It clarifies the structure of quantum $\alpha$-Rényi type divergences that satisfy the monotonicity under a subclass of quantum channels consisting of quantum-classical and classical-quantum channels. At the moment we have no explicit example of quantum $\alpha$-Rényi divergence that is not monotone under all quantum channels but monotone under the above subclass. Although Appendix C has not been used in the main body of this paper, it might be of independent interest from the point of view of quantum information.
(5) It is well known [18, 19, 35, 37] that the operator perspective
$\mathcal{P}_{f}(B,A):=A^{1/2}f(A^{-1/2}BA^{-1/2})A^{1/2}$
associated to a function $f$ on $(0,\infty)$ is jointly operator convex on $\mathbb{M}_{n}^{++}\times\mathbb{M}_{n}^{++}$, $n\in\mathbb{N}$, if $f$ is operator convex, while it is jointly operator concave if $f$ is operator monotone. Thus it might be expected to obtain a general joint convexity theorem for $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}$ for certain $p>0$, that is the convexity counterpart of Theorem 5.1 when $f$ is operator convex on $(0,\infty)$. However, this is not possible except for $p=1$, in the case that $\mathcal{P}_{f}$ is jointly operator convex. Indeed, when $f$ is a linear function $f(x)=1-\alpha+\alpha x$ ($0<\alpha<1$), the joint convexity of $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}=\mathrm{Tr}\,\mathcal{A}_{\alpha,p}(A,B)$ restricts to $1\leq p\leq 2$ by Proposition 5.2(2). When $f(x)=x^{\alpha}$ ($1<\alpha\leq 2$), the joint convexity of $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}=\mathrm{Tr}\,G_{\alpha,p}$ restricts to $p\leq 1$. The intersection is only $p=1$. It might be possible that $\mathrm{Tr}\,\mathcal{P}_{f}(B^{p},A^{p})^{1/p}$ is jointly convex for a certain $p<1$ under the constraint of $f(0^{+})=0$.
Acknowledgements
The author thanks Milan Mosonyi for invitation to the workshop at the Erdős Center in July, 2024 and for valuable suggestions which helped to improve this paper.
Appendix A Equality case in norm inequality for $G_{\alpha,p}$
Theorem A.1.
Let $\alpha\in(0,2]\setminus\{1\}$ and $\|\cdot\|$ be a strictly increasing unitarily invariant norm on $\mathbb{M}_{n}$. Then for every $A,B\in\mathbb{M}_{n}^{++}$ the following conditions are equivalent:
  • (i)
    $\|G_{\alpha,p}(A,B)\|=\|G_{\alpha,q}(A,B)\|$ for some $0<p<q$;
  • (ii)
    $\|G_{\alpha,p}(A,B)\|=\|LE_{\alpha}(A,B)\|$ for all $p>0$;
  • (iii)
    $G_{\alpha,p}(A,B)=G_{\alpha,q}(A,B)$ for some $0<p<q$;
  • (iv)
    $G_{\alpha,p}(A,B)=LE_{\alpha}(A,B)$ for all $p>0$;
  • (v)
    $AB=BA$.
Proof.
[…]
$\mathrm{Tr}\,G_{\alpha}(A_{0},B_{0})=\mathrm{Tr}\,G_{\alpha}(A_{0}^{t},B_{0}^{t})^{1/t},\qquad 1\leq t\leq q/p.$ (A.1)
Since the function $t\mapsto G_{\alpha}(A_{0}^{t},B_{0}^{t})^{1/t}$ is analytic in $(0,\infty)$, it follows that (A.1) extends to all $t\in(0,\infty)$. Now write $A_{0}=e^{H}$ and $B_{0}=e^{K}$ with Hermitian $H,K\in\mathbb{M}_{n}$. Since $G_{\alpha}(A_{0}^{t},B_{0}^{t})^{1/t}\to e^{\alpha H+(1-\alpha)K}$ as $t\to 0$, we obtain
$\mathrm{Tr}\,G_{\alpha}(e^{tH},e^{tK})^{1/t}=\mathrm{Tr}\,e^{\alpha H+(1-\alpha)K},\qquad t>0.$
By [39, Theorem 3.1] and [28, Lemma 2.2] again this implies that $\lambda\bigl(G_{\alpha}(e^{tH}t,e^{tK})^{1/t}\bigr)=\lambda(e^{\alpha H+(1-\alpha)K})$ for all $t>0$, so that
$\mathrm{Tr}\,G_{\alpha}(e^{tH},e^{tK})=\mathrm{Tr}\,e^{t(\alpha H+(1-\alpha)K)},\qquad t\in(-\infty,\infty).$ (A.2)
Below we compute the coefficients up to the $4$th of the Taylor expansion in $t$ of the LHS of (A.2). From the Taylor expansions of $e^{tH}$ and $e^{-tK/2}$ one can write
$e^{-tK/2}e^{tH}e^{-tK/2}=I+tX_{1}+t^{2}X_{2}+t^{3}X_{3}+t^{4}X_{4}+\cdots,$
where
$X_{1}$ $=H-K,$ (A.3)
$X_{2}$ $={(H-K)^{2}\over 2},$
$X_{3}$ $={H^{3}\over 6}-{H^{2}K+KH^{2}\over 4}+{HK^{2}+2KHK+K^{2}H\over 8}-{K^{3}\over 6},$
$X_{4}$ $={H^{4}\over 24}-{H^{3}K+KH^{3}\over 12}+{H^{2}K^{2}+2KH^{2}K+K^{2}H^{2}\over 16}$
$\qquad-{HK^{3}+3KHK^{2}+3K^{2}HK+K^{3}H\over 48}+{K^{4}\over 24}.$
Using
$(1+x)^{\alpha}=1+\alpha x+{\alpha(\alpha-1)\over 2}x^{2}+{\alpha(\alpha-1)(\alpha-2)\over 6}x^{3}+{\alpha(\alpha-1)(\alpha-2)(\alpha-3)\over 24}x^{4}+\cdots,$
one can write
$(e^{-tK/2}e^{tH}e^{-tK/2})^{\alpha}=I+tY_{1}+t^{2}Y^{2}+t^{3}Y_{3}+t^{4}Y_{4}+\cdots,$
where
$Y_{1}$ $=\alpha X_{1},$ (A.4)
$Y_{2}$ $=\alpha X_{2}+{\alpha(\alpha-1)\over 2}X_{1}^{2},$
$Y_{3}$ $=\alpha X_{3}+{\alpha(\alpha-1)\over 2}(X_{1}X_{2}+X_{2}X_{1})+{\alpha(\alpha-1)(\alpha-2)\over 6}X_{1}^{3},$
$Y_{4}$ $=\alpha X_{4}+{\alpha(\alpha-1)\over 2}(X_{1}X_{3}+X_{2}^{2}+X_{3}X_{1})$
$\qquad+{\alpha(\alpha-1)(\alpha-2)\over 6}(X_{1}^{2}X_{2}+X_{1}X_{2}X_{1}+X_{2}X_{1}^{2})+{\alpha(\alpha-1)(\alpha-2)(\alpha-3)\over 24}X_{1}^{4}.$
Therefore, one can write
$\mathrm{Tr}\,G_{\alpha}(e^{tH},e^{tK})=\mathrm{Tr}(e^{tK/2}e^{tH}e^{-tK/2})^{\alpha}e^{tK}=\mathrm{Tr}\,I+tz_{1}+t^{2}z_{2}+t^{3}z_{3}+t^{4}z_{4}+\cdots,$
where
$z_{1}$ $=\mathrm{Tr}(Y_{1}+K),$ (A.5)
$z_{2}$ $=\mathrm{Tr}\biggl(Y_{2}+Y_{1}K+{K^{2}\over 2}\biggr),$
$z_{3}$ $=\mathrm{Tr}\biggl(Y_{3}+Y_{2}K+Y_{1}{K^{2}\over 2}+{K^{3}\over 6}\biggr),$
$z_{4}$ $=\mathrm{Tr}\biggl(Y_{4}+Y_{3}K+Y_{2}{K^{2}\over 2}+Y_{1}{K^{3}\over 6}+{K^{4}\over 24}\biggr).$
Direct though quite tedious computations from (A.3)–(A.5) give
$z_{1}$ $:=\mathrm{Tr}(\alpha H+(1-\alpha)K),$
$z_{2}$ $:={1\over 2}\mathrm{Tr}(\alpha H+(1-\alpha)K)^{2},$
$z_{3}$ $:={1\over 6}\mathrm{Tr}(\alpha H+(1-\alpha)K)^{3},$
$z_{4}$ $:={1\over 24}\mathrm{Tr}\Bigl(\alpha^{4}H^{4}+4\alpha^{3}(1-\alpha)H^{3}K+4\alpha(\alpha-1)(\alpha^{2}-\alpha+1)H^{2}K^{2}$
$\qquad\qquad+2\alpha(\alpha-1)(\alpha^{2}-\alpha-2)HKHK+4\alpha(1-\alpha)^{3}HK^{3}+(1-\alpha)^{4}K^{4}\Bigr).$ (A.6)
Comparing the Taylor coefficients of $t^{4}$ of both sides of (A.2), we have
$24z_{4}$ $=\mathrm{Tr}(\alpha H+(1-\alpha)K)^{4}$
$=\mathrm{Tr}\Bigl(\alpha^{4}H^{4}+4\alpha^{3}(1-\alpha)H^{3}K+4\alpha^{2}(1-\alpha)^{2}H^{2}K^{2}$
$\qquad\quad+2\alpha^{2}(1-\alpha)^{2}HKHK+4\alpha(1-\alpha)^{3}HK^{3}+(1-\alpha)^{4}K^{4}\Bigr).$
From this and (A.6) we finally arrive at
$\mathrm{Tr}\,H^{2}K^{2}=\mathrm{Tr}\,HKHK,$
which is equivalently written as $\mathrm{Tr}(HK-KH)^{*}(HK-KH)=0$. Hence $HK=KH$, so that $A_{0}B_{0}=B_{0}A_{0}$, equivalently $AB=BA$. ∎
Appendix B Proof of Theorem 5.3
Proof.
[…]
$\Phi\biggl(\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}\biggr):=X_{11}+X_{22},\qquad X_{ij}\in\mathbb{M}_{n}.$
For any $A_{i},B_{i}\in\mathrm{dom}_{n}\,Q$, $i=1,2$, and $\lambda\in(0,1)$ set $A:=\lambda A_{1}\oplus(1-\lambda)A_{2}$ and $B:=\lambda B_{1}\oplus(1-\lambda)B_{2}$ in $\mathbb{M}_{2n}$. We then have
$Q(\lambda A_{1}+(1-\lambda)A_{2},\lambda B_{1}+(1-\lambda)B_{2})$
$\quad=Q(\Phi(A),\Phi(B))$
$\quad\geq Q(\lambda A_{1}\oplus(1-\lambda)A_{2},\lambda B_{1}\oplus(1-\lambda)B_{2})\qquad\mbox{(by (i))}$
$\quad=\lambda Q(A_{1},B_{1})+(1-\lambda)Q(A_{2},B_{2})\qquad\mbox{(by (c), (b))}.$
[…]
$\Phi(X)=\mathrm{Tr}_{nl}V(X\otimes\eta)V^{*},\qquad X\in\mathbb{M}_{n},$ (B.1)
where $\mathrm{Tr}_{nl}:\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}\to\mathbb{M}_{m}$ is the partial trace. Let $\tau_{0}:=(nl)^{-1}I_{nl}\in\mathbb{M}_{n}\otimes\mathbb{M}_{l}$. Then $\tau_{0}\otimes\mathrm{Tr}_{nl}$ is the conditional expectation from $\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}$ onto $I_{nl}\otimes\mathbb{M}_{m}$ with respect to the trace, and it is well known (see, e.g., [51, Sec. 7]) that for every $Z\in\mathbb{M}_{n}\otimes\mathbb{M}_{l}\otimes\mathbb{M}_{m}$,
$\tau_{0}\otimes\mathrm{Tr}_{nl}(Z)=(nl)^{-2}\sum_{\mu,\nu=1}^{nl}(S^{\mu}\otimes I_{m})(W^{\nu}\otimes I_{m})Z(W^{*\mu}\otimes I_{m})(S^{*\nu}\otimes I_{m}),$ (B.2)
where $S$ and $W$ are the Weyl–Heisenberg matrices in $\mathbb{M}_{d}$, $d:=nl$, i.e., $S_{jk}:=\delta_{j+1,k}$ (with $j+1$ mod $d$) and $W_{jk}:=\omega^{j}\delta_{jk}$ with $\omega:=e^{2\pi i/d}$. Note that $U_{\mu\nu}:=(S^{\mu}\otimes I_{m})(W^{\nu}\otimes I_{m})$ is a unitary and $U_{\mu\nu}^{*}=(W^{*\mu}\otimes I_{m})(S^{*\nu}\otimes I_{m})$. By (B.1) and (B.2) we have
$\tau_{0}\otimes\Phi(X)=(nl)^{-2}\sum_{\mu,\nu=1}^{nl}U_{\mu\nu}V(X\otimes\eta)V^{*}U_{\mu\nu}^{*},\qquad X\in\mathbb{M}_{n}.$ (B.3)
Therefore,
$Q(\Phi(A),\Phi(B))$ $=Q(\tau_{0}\otimes\Phi(A),\tau_{0}\otimes\Phi(B))\qquad\mbox{(by (d), (a))}$
$=Q\Biggl((nl)^{-2}\sum_{\mu,\nu=1}^{nl}U_{\mu\nu}V(A\otimes\eta)V^{*}U_{\mu\nu}^{*},(nl)^{-2}\sum_{\mu,\nu=1}^{nl}U_{\mu\nu}V(B\otimes\eta)V^{*}U_{\mu\nu}^{*}\Biggr)$
$\geq(nl)^{-2}\sum_{\mu,\nu=1}^{nl}Q\bigl(U_{\mu\nu}V(A\otimes\eta)V^{*}U_{\mu\nu}^{*},U_{\mu\nu}V(B\otimes\eta)V^{*}U_{\mu\nu}^{*}\bigr)\qquad\mbox{(by (ii))}$
$=Q(A\otimes\eta,B\otimes\eta)\qquad\mbox{(by (e))}$
$=Q(A,B)\qquad\mbox{(by (d), (a))}.$
[…]
Appendix C Monotonicity under restricted CPTP maps
Let $\alpha\in(0,\infty)\setminus\{1\}$ and $Q_{\alpha}:(A,B)\in\mathrm{dom}\,Q_{\alpha}\mapsto Q_{\alpha}(A,B)\in[0,+\infty)$ be a function with the domain $\mathrm{dom}\,Q_{\alpha}$ given in (5.1), where $\mathrm{dom}\,Q_{\alpha}=\bigsqcup_{n}(\mathbb{M}_{n}^{+}\times\mathbb{M}_{n}^{+})_{\geq}$ when $\alpha>1$. Assume that $Q_{\alpha}$ satisfies properties (d) and (e) in Section 5.1 and
$Q_{\alpha}(\mathrm{diag}(a),\mathrm{diag}(b))=\sum_{i=1}^{n}a_{i}^{1-\alpha}b_{i}^{\alpha}$ (C.1)
for all $a,b\in[0,\infty)^{n}$, $n\in\mathbb{N}$, such that $(\mathrm{diag}(a),\mathrm{diag}(b))\in\mathrm{dom}\,Q_{\alpha}$. Define the quantum divergence of Renyi type defined for $A,B\in\mathbb{M}_{n}^{+}$, $n\in\mathbb{N}$, with $B\neq 0$ as in (5.7) by
$D^{Q_{\alpha}}(B\|A):=\begin{cases}{1\over\alpha-1}\log{Q_{\alpha}(A,B)\over\mathrm{Tr}\,B}&\text{if $(A,B)\in\mathrm{dom}Q_{\alpha}$},\\ +\infty.&\text{otherwise}\end{cases}$ (C.2)
For example, if $\mathcal{M}_{\alpha,p}$ is any of $R_{\alpha,p}$, $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$ for $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$, then $Q_{\alpha}:=\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ meets the above requirements. In Theorem 5.8 four trace inequalities have appeared as necessary conditions for $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ to be jointly concave or jointly convex. In the next theorem we show what is meant by each of those inequalities for more general $Q_{\alpha}$ stated above.
We consider two special classes of quantum channels (CPTP maps). If $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$ is a positive trace-preserving map such that the range of $\Phi$ is included in a commutative subalgebra of $\mathbb{M}_{m}$ is called a quantum-classical channel. In this case, there are orthogonal projections $(P_{i})_{i=1}^{k}$ in $\mathbb{M}_{m}$ such that $\sum_{i=1}^{k}P_{i}=I_{m}$ and the range of $\Phi$ is included in $\bigoplus_{i=1}^{k}\mathbb{C}P_{i}$. Then we have a POVM $(M_{i})_{i=1}^{k}$ on $\mathbb{C}^{n}$ (see (5.3)) such that $\Phi(A)=\sum_{i=1}^{k}(\mathrm{Tr}\,M_{i}A/\mathrm{Tr}\,P_{i})P_{i}$ for $A\in\mathbb{M}_{n}$. In this way, a quantum-classical channel is essentially a POVM. The other way around, a positive trace preserving map $\Phi:\mathbb{C}^{n}\to\mathbb{M}_{m}$ is called a classical-quantum channel. Those quantum-classical and classical-quantum channels are are automatically CPTP maps.
Theorem C.1.
Let $Q_{\alpha}$ be as mentioned above.
  • (1)
[…]
    • (i)
      $Q_{\alpha}(A,B)\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$ for all $(A,B)\in\mathrm{dom}\,Q_{\alpha}$.
    • (ii)
      $Q_{\alpha}(\Phi(A),\Phi(B))\geq Q_{\alpha}(A,B)$ for all $(A,B)\in\mathrm{dom}_{n}Q_{\alpha}$ and for any quantum-classical channel $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{m}$, $n,m\in\mathbb{N}$.
    • (iii)
      $Q_{\alpha}(A,\mathcal{E}_{A}(B))\geq Q_{\alpha}(A,B)$ for all $(A,B)\in\mathrm{dom}\,Q_{\alpha}$, where $\mathcal{E}_{A}$ is the pinching with respect to $A$.
  • (2)
    Let $1<\alpha<\infty$. The following conditions (i)–(iii) are equivalent:
    • (i)
      $Q_{\alpha}(A,B)\geq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$ for all $(A,B)\in\mathrm{dom}\,Q_{\alpha}$.
    • (ii)
      $Q_{\alpha}(\Phi(A),\Phi(B))\leq Q_{\alpha}(A,B)$ for all $(A,B)\in\mathrm{dom}_{n}Q_{\alpha}$ and for any quantum-classical channel $\Phi:\mathbb{M}_{n}\to\mathbb{M}_{n}$, $n,m\in\mathbb{N}$.
    • (iii)
      $Q_{\alpha}(A,\mathcal{E}_{A}(B))\leq Q_{\alpha}(A,B)$ for all $(A,B)\in\mathrm{dom}\,Q_{\alpha}$.
  • (3)
    Let $0<\alpha<1$. The following conditions (i) and (ii) are equivalent:
    • (i)
      $Q_{\alpha}(A,B)\geq\mathrm{Tr}\,G_{\alpha,1}(A,B)$ for all $(A,B)\in\mathrm{dom}\,Q_{\alpha}$.
    • (ii)
      $Q_{\alpha}(\Phi(a),\Phi(b))\geq Q_{\alpha}(\mathrm{diag}(a),\mathrm{diag}(b))$ for all $a,b\in[0,\infty)^{n}$ with $(\mathrm{diag}(a),\mathrm{diag}(b))\in\mathrm{dom}\,Q_{\alpha}$ and for any classical-quantum channel $\Phi:\mathbb{C}^{n}\to\mathbb{M}_{m}$, $n,m\in\mathbb{N}$.
  • (4)
[…]
    • (i)
      $Q_{\alpha}(A,B)\leq\mathrm{Tr}\,G_{\alpha,1}(A,B)$ for all $(A,B)\in\mathrm{dom}\,Q_{\alpha}$.
    • (ii)
      $Q_{\alpha}(\Phi(a),\Phi(b))\leq Q_{\alpha}(\mathrm{diag}(a),\mathrm{diag}(b))$ for all $a,b\in[0,\infty)^{n}$ with $(\mathrm{diag}(a),\mathrm{diag}(b))\in\mathrm{dom}\,Q_{\alpha}$ and for any classical-quantum channel $\Phi:\mathbb{C}^{n}\to\mathbb{M}_{m}$, $n,m\in\mathbb{N}$.
Proof.
We give the proofs of (1) and (4) only; those of (2) and (3) are similar to (1) and (4), respectively. First note that $Q_{\alpha}(A,0)=0$ for all $A\geq 0$. In fact, by property (e) and (C.1) we have $Q_{\alpha}(A,0)=Q_{\alpha}(\mathrm{diag}(\lambda(A)),\mathrm{diag}(0))=0$, where $\lambda(A)$ is the eigenvalue vector of $A$. Since $R_{\alpha,1/\alpha}(A,0)=G_{\alpha,1}(A,0)=0$ as well, we may always assume $B\neq 0$ and $b\neq 0$ in the following.
[…]
$D^{Q_{\alpha}}(\Phi(B)\|\Phi(A))=D_{\alpha}^{\mathrm{cl}}(\Phi(B)\|\Phi(A))=\widetilde{D}_{\alpha}(\Phi(B)\|\Phi(A))\leq\widetilde{D}_{\alpha}(B\|A)\leq D^{Q_{\alpha}}(B\|A).$
This implies (ii).
[…]
$Q_{\alpha}(A,\mathcal{E}_{A}(B))=Q_{\alpha}(\mathrm{diag}(a),\mathrm{diag}(b))=Q_{\alpha}(\Phi(A),\Phi(B))\geq Q_{\alpha}(A,B).$
[…]
$D^{Q_{\alpha}}(B\|A)={1\over m}D^{Q_{\alpha}}(B^{\otimes m}\|A^{\otimes m})\geq{1\over m}D^{Q_{\alpha}}(\mathcal{E}_{A^{\otimes m}}(B^{\otimes m})\|A^{\otimes m})={1\over m}D_{\alpha}^{\mathrm{cl}}(\mathcal{E}_{A^{\otimes m}}(B^{\otimes m})\|A^{\otimes m})$
where the first equality is by definition (C.2) and property (d), the inequality is due to (iii), and the last equality is by (e) and (C.1). Since $1/2\leq\alpha<1$, letting $m\to\infty$ (see definition (5.4)) and applying Theorem 5.7(3) we have $D^{Q_{\alpha}}(B\|A)\geq\widetilde{D}_{\alpha}(B\|A)=D^{R_{\alpha,1/\alpha}}(B\|A)$, which implies (i).
[…]
$D^{Q_{\alpha}}(\Phi(b)\|\Phi(a))\leq D^{G_{\alpha,1}}(\Phi(b)\|\Phi(a))\geq D_{\alpha}^{\max}(\Phi(b)\|\Phi(a)).$
Moreover, we have $D_{\alpha}^{\max}(\Phi(b)\|\Phi(a))\leq D_{\alpha}^{\mathrm{cl}}(b\|a)=D^{Q_{\alpha}}(\mathrm{diag}(b)\|\mathrm{diag}(a))$ by (5.2) and (C.1), so that $D^{Q_{\alpha}}(\Phi(b)\|\Phi(a))\leq D^{Q_{\alpha}}(\mathrm{diag}(b)\|\mathrm{diag}(a))$, implying (ii).
[…]
$D^{Q_{\alpha}}(B\|A)=D^{Q_{\alpha}}(\Gamma(b)\|\Gamma(a))\leq D^{Q_{\alpha}}(\mathrm{diag}(b)\|\mathrm{diag}(a))=D_{\alpha}^{\mathrm{cl}}(b\|a).$
Therefore, it follows from (5.2) and Theorem 5.7(1) that $D^{Q_{\alpha}}(B\|A)\leq D_{\alpha}^{\max}(B\|A)=D^{G_{\alpha,1}}(B\|A)$, implying (i). ∎
Example C.2.
Here let us exemplify Theorem C.1 for $Q_{\alpha}=\mathrm{Tr}\,R_{\alpha,p}$ and $\mathrm{Tr}\,G_{\alpha,p}$. We say that a quantum channel $\Phi$ is semi-classical if it is either quantum-classical or classical-quantum.
(1) As for $\mathrm{Tr}\,R_{\alpha,p}$, by Theorems 3.1(a) and 3.2 note that for $0<\alpha<1$, $\mathrm{Tr}\,G_{\alpha,1}\leq\mathrm{Tr}\,R_{\alpha,p}\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}$ if and only if $p\leq 1/\alpha$, and that for $\alpha>1$, $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,R_{\alpha,p}\leq\mathrm{Tr}\,G_{\alpha,1}$ if and only if $\max\{\alpha/2,\alpha-1\}\leq 1/p\leq\alpha$. Theorem C.1 says that if either of these holds, then $D^{R_{\alpha,p}}$ ($D_{\alpha,z}$ with $z=1/p$) is monotone under all semi-classical channels, and vice versa if $\alpha\geq 1/2$. Thus, in view of Theorem 5.9 we see that for any $\alpha\in[1/2,\infty)\setminus\{1\}$ the $\alpha$-$z$-Rényi divergence $D_{\alpha,z}$ satisfies the monotonicity under quantum channels if and only if it does under semi-classical channels.
(2) As for $\mathrm{Tr}\,G_{\alpha,p}$, by Proposition 4.2 and Theorem 3.2 note that for $0<\alpha<1$, $\mathrm{Tr}\,G_{\alpha,1}\leq\mathrm{Tr}\,G_{\alpha,p}\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}$ if and only if $p\leq 1$, and that for $1\leq\alpha\leq 2$, $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,G_{\alpha,p}\leq\mathrm{Tr}\,G_{\alpha,1}$ if and only if $\max\bigl\{1/2,{\alpha-1\over\alpha}\bigr\}\leq p\leq 1$. Theorem C.1 says that either of these holds if and only if $D^{G_{\alpha,p}}$ satisfies the monotonicity under semi-classical channels. Of course, this is only if $D^{G_{\alpha,p}}$ satisfies the monotonicity under quantum channels; see Theorem 5.15 and Proposition 5.16.
The next corollary shows the ‘conditional’ joint concavity/convexity of the above $Q_{\alpha}$.
Corollary C.3.
Let $Q_{\alpha}$ be as above and assume that it satisfies, in addition to (d), (e), properties (b), (c) in Section 5.1. Let $(A_{i},B_{i})\in\mathrm{dom}\,Q_{\alpha}$, $i=1,2$, and $0<\lambda<1$. Assume that $\lambda A_{1}+(1-\lambda)A_{2}$ and $\lambda B_{1}+(1-\lambda)B_{2}$ commute. If $0\leq\alpha<1$ and condition (ii) (or (i)) of Theorem C.1(1) holds, then
$Q_{\alpha}(\lambda A_{1}+(1-\lambda)A_{2},\lambda B_{1}+(1-\lambda)B_{2})\geq\lambda Q_{\alpha}(A_{1},B_{1})+(1-\lambda)Q_{\alpha}(A_{2},B_{2}).$
If $1<\alpha<\infty$ and condition (ii) (or (i)) of Theorem C.1(2) holds, then
$Q_{\alpha}(\lambda A_{1}+(1-\lambda)A_{2},\lambda B_{1}+(1-\lambda)B_{2})\leq\lambda Q_{\alpha}(A_{1},B_{1})+(1-\lambda)Q_{\alpha}(A_{2},B_{2}).$
Proof.
Let $A_{i},B_{i}$ be in $\mathbb{M}_{n}$. By assumption we take the commutative subalgebra $\mathcal{C}$ of $\mathbb{M}_{n}$ generated by $\lambda A_{i}+(1-\lambda)B_{i}$, $i=1,2$. Let $\mathcal{E}_{\mathcal{C}}$ be the trace-preserving conditional expectation from $\mathbb{M}_{n}$ onto $\mathcal{C}$. Define a quantum-classical channel $\Phi_{2n}:\mathbb{M}_{2n}=\mathbb{M}_{n}\otimes\mathbb{M}_{2}\to\mathbb{M}_{n}$ by
$\Phi\biggl(\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}\biggr):=\mathcal{E}_{\mathcal{C}}(X_{11}+X_{22}),\qquad X_{ij}\in\mathbb{M}_{n}.$
[…]
For instance, if $\alpha>1$ and $p\geq\max\bigl\{1/2,{\alpha-1\over\alpha}\}$, then Theorem 3.2(2) implies that $\mathrm{Tr}\,G_{\alpha,p}(A,B)\geq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$. Hence in this case, $\mathrm{Tr}\,G_{\alpha,p}$ satisfies the ‘conditional’ joint convexity as in the above corollary, supplementing Proposition 5.16(1).